### v1

# Graph-based Retrieval Augmented Generation for Dynamic Few-shot Text Classification

Abstract.
Text classification is a fundamental task in natural language processing, pivotal to various applications such as query optimization, data integration, and schema matching. While neural network-based models, such as CNN and BERT, have demonstrated remarkable performance in text classification, their effectiveness heavily relies on abundant labeled training data. This dependency makes these models less effective in dynamic few-shot text classification, where labeled data is scarce, and target labels frequently evolve based on application needs. Recently, large language models (LLMs) have shown promise due to their extensive pretraining and contextual understanding. Current approaches provide LLMs with text inputs, candidate labels, and additional side information (e.g., descriptions) to predict text labels. However, their effectiveness is hindered by the increased input size and the noise introduced through side information processing. To address these limitations, we propose a graph-based online retrieval-augmented generation framework, namely GORAG, for dynamic few-shot text classification. GORAG constructs and maintains an adaptive information graph by extracting side information across all target texts, rather than treating each input independently. It employs a weighted edge mechanism to prioritize the importance and reliability of extracted information and dynamically retrieves relevant context using a minimum-cost spanning tree tailored for each text input. Empirical evaluations demonstrate that GORAG outperforms existing approaches by providing more comprehensive and accurate contextual information.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at ***.
1. Introduction
Text classification is a fundamental task that is connected to various other tasks, such as query optimization (Yu et al., 2023; Yu and Litchfield, 2020) data integration (Dong and Rekatsinas, 2018), and schema matching (Sahay et al., 2020; Zhang et al., 2024a). For example, in schema matching tasks, a text classifier can be trained to determine whether two schema attributes refer to the same concept based on their descriptions or metadata (Zhang et al., 2024a). In recent years, Neural Network (NN)-based models (Johnson and Zhang, 2015; Prieto et al., 2016; Liu et al., 2016; Dieng et al., 2017; Wang et al., 2017; Johnson and Zhang, 2017; Peters et al., 2018; Wang et al., 2018; Yang et al., 2019; Miao et al., 2021), such as CNN (Kim, 2014), Bert (Devlin, 2018) and RoBERTa (Liu, 2019), have demonstrated impressive performance on text classification tasks. However, the effectiveness of these NN-based approaches depend on abundant labeled training data, which requires significant time and human efforts (Meng et al., 2018). Consequently, these methods perform poorly with limited labeled text data (Xia et al., 2021), such as in few-shot text classification tasks (Wang et al., 2020; Aljehani et al., 2024).
Figure 1. An illustration of the structural composition of current Reteieval Augmented Generation based few-shot text classification models, where the few-shot labeled training data is indexed as an external information source. During classification, the model retrieves side information from the external information source to provide additional context that aids the classification process.
Also, in real-world applications, such as Web Of Science (Kowsari et al., 2017) and IMDb (imd, [n.d.]), target labels for text often change based on the application’s requirements (Xia et al., 2021), leading to dynamic text classification. For instance, movie labels might initially be {Action, Thriller, Comedy}, but platforms may need to add more genres (e.g., Romantic and Comic) to extend the original labels. Therefore, it is important to classify text into these new categories. Consequently, how to develop methods to dynamically classify text with limited labeled data (i.e., dynamic few-shot text classification) remains a necessity and open problem.
Depending the technique for dynamic few-shot text classification task, current models can be categorized into three types, i.e., data augmentation-based (Meng et al., 2018; Meng et al., 2020; Xia et al., 2021), meta learning-based (Bansal et al., 2020; Zhang et al., 2023a), and large language model (LLM)-based approaches (Chen et al., 2023; Sarthi et al., 2024; Edge et al., 2024; Gutiérrez et al., 2024; Guo et al., 2024). Firstly, data augmentation-based models create additional training data by mixing the pairs of the few-shot labeled data and assign a mixed labels for these created data based on the labels of each pairs of data (Xia et al., 2021). Also, several data augmentation approaches create extra semantic related content based on the label names (Meng et al., 2018; Meng et al., 2020). However, due to the limited labeled data, the generated text data has very limited patterns. Consequently, the text classification models trained on these generated data are over-fitting on limited text data and are not generalizable (Liu et al., 2023). Also, meta-learning models (Bansal et al., 2020; Zhang et al., 2023a) train meta-learners on each base class to quickly adapt to new classes with minimal data. However, meta-learning still requires a substantial amount of labeled data for base classes.
Recently, large language models (Sun et al., 2021; Jiang et al., 2023; Touvron et al., 2023; Achiam et al., 2023; Yang et al., 2024), pretrained on extensive corpora, have achieved significant success due to their superior and comprehensive text understanding abilities. Several researchers (Chen et al., 2023; Sarthi et al., 2024) provide the text, candidate labels, and retrieve side information (e.g., descriptions of text and labels, or documents) for LLMs, and these LLMs output the labels of the text. Current work (Gao et al., 2023) demonstrates that this side information can provide context for the text and labels, enabling LLMs to comprehensively understand these inputs, which is a key component of the success of LLMs in text classification. However, the incorporation of side information can further increase the input size and noise (Chen et al., 2023), which impedes the efficiency and effectiveness of LLMs (Sarthi et al., 2024). As a result, compression-based retrieval-augmented generation (RAG) approaches (Edge et al., 2024; Gutiérrez et al., 2024; Guo et al., 2024) are proposed to construct graphs by extracting key information from the side information sources. These models then only retrieve the information (e.g., graph path or subgraphs) relevant to the target text as the side information for LLMs, reducing the length of inputs for LLMs.
However, existing compression-based RAG approaches (Edge et al., 2024; Guo et al., 2024; Gutiérrez et al., 2024) still have three limitations in the dynamic few-shot text classification task. Firstly, these approaches only extract information and build an information graph based on the side information of each text individually. In this way, they do not merge the information graphs of different text inputs into a single graph to obtain more comprehensive information for incoming target texts. Consequently, they fail to consider the correlation between side information of different text inputs, providing insufficient and incomplete context for LLMs. Secondly, existing approaches construct graphs by indexing extracted information linkages uniformly. They do not consider the varying importance and extraction confidence of each link, which may provide incorrect and unreliable context for LLMs. Thirdly, existing approaches select relevant information for each input text based on a globally predefined threshold. However, the optimal retrieval threshold can vary across different text samples, making the globally predefined threshold suboptimal for the entire dataset.
To address the aforementioned issues, we propose a novel Graph-based Online Retrieval Augmented Generation framework for dynamic few-shot text classification, called GORAG. In general, GORAG constructs and maintains an adaptive online information graph by extracting side information from all target texts, tailoring graph retrieval for each input. Firstly, GORAG extracts keywords from the text using the LLM and links these keywords with the text’s ground truth label to represent the relationship between keywords and labels. Then, GORAG employs an edge-weighting mechanism when indexing edges into the information graph, assigning edge weights based on the importance of the keywords and their relevance to the respective text’s label. Secondly, when retrieving information from the graph for each text, GORAG retrieves candidate labels by constructing a minimum-cost spanning tree of text keywords mapped to the graph. After retrieval, the candidate labels serve as a filtered subset of the original target labels, which are then used to create input for the LLM. Since the generated spanning tree is determined solely by the graph information and the extracted text keywords, GORAG achieves adaptive retrieval without relying on any human-defined retrieval thresholds.
We summarize the novel contributions as follows.
  • We present a RAG framework for few-shot text classification tasks, namely GORAG. The GORAG framework consists of four stages: (1) Graph construction, where a graph is built with keywords from the training texts considering their correlation with the texts’ label; (2) Candidate label retrieval, where relevant labels are adaptively retrieved from the graph for label pre-filtering; (3) Prompt construction, where prompts are generated based on the filtered candidate labels and the testing texts; and (4) Online indexing, where the keywords of the un-labeled text are indexed to the constructed graph based on the prediction result.
  • We develop a novel graph edge weighting mechanism based on the keyword nodes’ importance within the text corpus, which can be applied during both graph construction and online indexing. This mechanism enables our approach to effectively model the relevance between keywords and labels, thereby reducing the noise introduced to the graph.
  • To avoid the need for human-selected thresholds during retrieval, we formulate the candidate label retrieval problem which akin to the NP-hard Steiner Tree problem. To solve this problem, we modified the Mehlhorn algorithm (Mehlhorn, 1988) which provides an efficient and effective solution to the problem to generate candidate types from the information graph.
In the rest of this paper, we first present prelimnary and related work in section 2, the pipeline of our model GORAG in section 3, our experiments in section 4 and the final conclusion in section 5.
2. Preliminary and Related Works
In this section, we first introduce the preliminaries of dynamic few-shot text classification in Section 2.1 and then discuss the existing approaches in Section 2.2. The important notations used in this paper are listed in Table 1.
2.1. Dynamic Few-shot Text Classification
Text classification (Kowsari et al., 2019; Minaee et al., 2021; Li et al., 2022) is a key task in real-world application that involves assigning predefined labels $y\in\mathcal{Y}$ to text $t=(w_{1},\cdots,w_{|t|})$ based on its words $w_{i}$. It is widely applied in areas like sentiment analysis (Kouadri et al., 2020; maamar2022sa),query optimization (Yu et al., 2023; Yu and Litchfield, 2020) data integration (Dong and Rekatsinas, 2018), and schema matching (Sahay et al., 2020; Zhang et al., 2024a). Traditional approaches rely on large labeled datasets (Devlin, 2018; Liu, 2019), which may not always be available. Recently, few-shot learning addresses this limitation by enabling models to classify text with only a small number of labeled examples per class (Meng et al., 2018; Meng et al., 2020). Dynamic classification (Parisi et al., 2019; Hanmo et al., 2024; Wang et al., 2024) introduces an additional challenge where new classes are introduced over multiple rounds, requiring the model to adapt to new classes while retaining knowledge of previously seen ones. Combining these aspects, Dynamic Few-Shot Text Classification (DFSTC) (Xia et al., 2021) allows the model to handle evolving classification tasks with minimal labeled data.
In DFSTC, the model is provided with multiple rounds of new class updates. Specifically, In each round $r$, a new set of classes $\mathcal{Y}_{new}^{r}$ is introduced and the labelled dataset for $\mathcal{Y}_{\text{new}}^{r}$ is denoted as $\mathcal{D}_{new}^{r}=\cup_{y_{i}\in\mathcal{Y}_{\text{new}}^{r}}\{t_{j},y_{i}\}_{j=1}^{k}$, where per class $y\in\mathcal{Y}_{\text{new}}^{r}$ only has $k$ labeled examples $\{t_{j},y_{i}\}_{j=1}^{k}$. Also, we denote candidate cumulative labels from the first round to the $r$-th round as $\mathcal{Y}^{r}=\bigcup_{i=1}^{r}\mathcal{Y}_{\text{new}}^{i}$. Formally, at the $r$-th round, given the candidate labels $\mathcal{Y}^{r}$ and all labeled data $\mathcal{D}^{r}=\cup_{i=1}^{r}\mathcal{D}_{new}^{i}$, the target of DFSTC task is to learn a function $f^{r}_{\theta}$, which can learn scores for all target labels $\mathbf{s}^{r}_{k}={f}^{r}_{\theta}(t,\mathcal{Y}^{r})\in[0,1]^{|\mathcal{Y}^{r}|}$ for the unseen text $t_{k}$. Then, we can get the predicted label $y^{r*}_{k}\in\mathcal{Y}^{r}$ for the unseen text $t_{k}$ as follows.
(1) $y^{r*}_{k}=\arg\max_{y^{r}_{k}\in\mathcal{Y}^{r}}({f}_{\theta}(t,\mathcal{Y}^{r})[y^{r}_{k}])$
DFSTC is valuable for its ability to learn from limited labeled data, adapt to evolving class distributions, and address real-world scenarios where categories and data evolve over time. For example, in Round 1, the model might be trained on classes like Sports, Politics, and Technology. In Round 2, new classes such as Health and Education are introduced. The model must now classify texts into all five classes while retaining its knowledge of earlier ones. For instance, given the text ”The new vaccine shows promising results,” the model should classify it as Health.
2.2. Dynamic Few-shot Text Classification Models
Current dynamic few-shot text classification models can be broadly categorized into two types: Meta learning-based models and Data Augmentation-based models, and Large Language Model (LLM)-based models.
2.2.1. Meta learning-based models
Meta learning-based models train a meta-learner on base classes $C_{b}$ and then adapt it to new classes in each round. For instance, LEOPARD (Bansal et al., 2020) use the meta-learner to generate model parameters to predict classes in each round. ConEntail (Zhang et al., 2023a) trains the meta-learner to classify texts based on their nested entailments with class label names at each round. However, these models require a sufficient amount of training data to train the meta-learner, which differs from our experimental setting and limits their generalization ability.
2.2.2. Data Augmentation-based models
Data Augmentation-based models generate additional data contrastively based on the few-shot labeled data to train the classifier. For example, Entailment (Xia et al., 2021) creates extra data samples, by concatenating texts with class label names. WeSTClass (Meng et al., 2018) trains a document generator to produce pseudo documents contrastively based on few-shot training data. LOTClass (Meng et al., 2020) uses a Pretrained Language Model (PLM) to generate words semantically correlated with the class label name to help with the classification. However, due to the limited labeled data, the generated text data of these models can have very limited patterns, which makes them prone to overfitting (Liu et al., 2023).
Table 1. Summary on Important Notations.
Notations Meanings
$t$ Text.
$w$ The edge weight.
$r$ The round of label updation.
$y\in\mathcal{Y}^{r}$ The target label at round $r$.
$k$ The number of labeled data provided per target label.
$f^{r}_{\theta}$ The function that learn all target labels’ score for texts.
$\mathbf{s}^{r}_{k}$ The score for all target labels at round $r$.
$y^{r*}_{k}$ The predicted label for the text.
$\mathcal{G}^{r}(\mathcal{V},\mathcal{E})$ Information graph with node set $\mathcal{V}$ and edge set $\mathcal{E}.$
$\mathcal{Y}^{r}$, $\mathcal{Y}^{r}_{new}$ All target label and new target labels at round $r$.
$\mathcal{D}^{r}$, $\mathcal{D}^{r}_{new}$ The labeled text for all labels and new labels at round $r$.
$\mathcal{T}^{r}$, $\mathcal{T}_{new}^{r}$ The text of all labels and new labels at round $r$.
$\mathcal{T}_{test}^{r}$ The testing text for at round $r$.
$\mathcal{K}_{new}^{r}$ The extracted keyword set for new labels of round $r$.
$\mathcal{V}_{new}^{r}$ The set of new graph nodes to be added in round $r$.
$\mathcal{N}(\cdot)$ The neighbor set of a node in the information graph.
$p_{extract}$ The extraction instruction prompt.
$p_{gen}$ The generation instruction prompt.
$p_{classify}$ The classification instruction prompt.
$CS(\cdot)$ The correlation score between keywords and labels.
$w^{r}_{v,y}$ The weight of edge $e_{v,y}$ in round $r$.
$P_{u,v}$ The shortest path between node $u$ and node $v$.
$\mathcal{V}_{exist}$ The keywords that exist in graph.
$\mathcal{V}_{notexist}$ The keywords do not exist in graph.
$\mathcal{E}_{oi}^{t}$ Edges added based on text $t$.
Figure 2. An overview of the GORAG model is as follows. In Part 1, GORAG constructs a graph based on keywords extracted from the few-shot training data, taking into account the keyword extraction confidence and relatedness to the target labels. In Part 2, GORAG performs adaptive graph retrieval and utilizes the retrieved graph information to filter the target labels. In Part 3, GORAG carry out classification with the testing texts and the candidate labels retrieved from the information graph. In Part 4, GORAG applies online indexing to update the graph with newly extracted nodes from testing texts through online indexing.
2.2.3. LLM-based models
Recently, Large Language Model (LLM)-based models have undergone rapid development (Fernandez et al., 2023; Amer-Yahia et al., 2023; Zhang et al., 2024b; Miao et al., 2024) and have been successfully adapted to various tasks, including data discovery (Arora et al., 2023; Dong et al., 2023; Kayali et al., 2024), entity or schema matching (Zhang et al., 2023b; Tu et al., 2023; Fan et al., 2024; Du et al., 2024), and natural language to SQL conversion (Li et al., 2024; Ren et al., 2024; Trummer, 2022; Gu et al., 2023). Notably, LLMs are inherently capable of inference without fine-tuning (Sun et al., 2021; Jiang et al., 2023; Touvron et al., 2023; Achiam et al., 2023; Yang et al., 2024), making them originally suitable for dynamic text classification tasks. However, the lack of fine-tuning can lead LLMs to generate incorrect answers, as they lack task-specific knowledge, these incorrect answers are often referred to as hallucinations (Zhang et al., 2023c). To mitigate hallucinations, researchers have provided LLMs with side information for classification, namely long context RAG models. For example, Propositionizer (Chen et al., 2023) applies a fine-tuned LLM to convert side information into atomic expressions, to facilitate fine-grained information retrieval. RAPTOR (Sarthi et al., 2024) clusters side information and then generate summaries for each cluster, to help with the model prediction. However, the retrieved contents from these models remain unstructured and can be lengthy, which impedes the efficiency and effectiveness of LLMs, leading to lost-in-the-middle issue (Liu et al., 2024).
To address this, compression-based RAG models were proposed, these models try to compress side information to reduce the input length. Based on how they compress model inputs, these models can be classified into prompt compressor models, and graph-based RAG models. On the one hand, prompt compressor models, such as LLMLingua2 (Pan et al., 2024), apply LLM’s generation perplexity to filter out un-important tokens in the model input. Based on LLMLingua2, LongLLMLingua (Jiang et al., 2024) further considers the instruction to the model when compression, make the model instruction-aware. On the other hand, graph-based RAG models index side informations into a graph, and retrieve graph components or summaries, which are shorter than long contexts thet traditional RAG models retrieved. For instance, GraphRAG (Edge et al., 2024) constructs a graph and aggregates nodes into graph communities, then generate community summaries to help the LLM prediction. However, GraphRAG requires re-creating the community after each graph index, which incurs significant computational overhead.
To address this limitation of GraphRAG, two alternative models, HippoRAG (Gutiérrez et al., 2024) and LightRAG (Guo et al., 2024), have been proposed, which skip the formulation of communities and directly retrieve nodes or pathes from graphs. HippoRAG (Gutiérrez et al., 2024) retrieve nodes whose retrieval score is above a human defined threshold. However, the optimal retrieval threshold can vary across different data samples, and a globally fixed threshold selected by human may be suboptimal for the entire dataset. To address this issue, LightRAG (Guo et al., 2024) calculates the embedding similarity of text extracted entities with graph nodes, to achieve a one-to-one mapping and retrieve all triples involved these graph nodes.
3. Methdology
To address the aforementioned issues, we propose GORAG, a novel approach that achieves adaptive retrieval by extracting valuable side information from a minimum-cost spanning tree generated on the constructed graph. Specifically, GORAG takes into account each keyword’s with in the training and texting text corpus and relatedness to the class label by assigning different costs to edges on the constructed graph, allowing for more informed retrieval decisions. Moreover, GORAG incorporates an online indexing mechanism, which enables the model to index valuable information extracted from the testing texts to the constructed graph in real-time, thereby enhancing its ability to make accurate predictions in the future.
3.1. Framework Overview
As illustrated in Figure 2, our model consists of four primary components, i.e., Graph Construction, Graph Retrieval, Classification, and Online Indexing.
Figure 3. An example of the constructed information graph from WOS dataset, where red nodes denote the label node and black nodes denote the keyword node. In this subgraph, node Entity and node Neighborhood are considered most related to the label Relational Database, the edge between them with the label have the lowest cost assigned by our edge weighting mechanism.
Part 1: Graph Construction Graph construction proposes is to construct or update the information graph $\mathcal{G}^{r}_{new}(\mathcal{V}^{r}_{new},\mathcal{E}^{r}_{new},\mathcal{W}^{r}_{new})$ at round $r$ based on the labeled data $\mathcal{D}_{new}^{r}=\cup_{y_{i}\in\mathcal{Y}_{\text{new}}^{r}}\{t_{j},y_{i}\}_{j=1}^{k}$ for at the $r$-th round. A small subgraph of the information graph created on WOS dataset is shown in Figure 3, graph formed like this will be used to provide retrieved-augmented information as context for unseen text, enabling LLMs to better understand the unseen text and accurately predict its labels. Specifically, under each new data $\mathcal{D}_{new}^{r}$ at the $r$-th round, we first extract the new keywords $\mathcal{K}_{new}^{r}\in\{t_{j}\}_{j=1}^{|\mathcal{D}_{new}^{r}|}$ from the few-shot training texts $\mathcal{T}^{r}_{new}=\{t_{j}\}_{j=1}^{|\mathcal{D}_{new}^{r}|}$. Then, we will then assign an edge $e^{r}_{v,y}\in\mathcal{E}^{r}_{new}$ between each keyword $v\in\mathcal{K}_{new}^{r}$ and each label $y\in\mathcal{Y}^{r}_{new}$. We then compute the weight $w^{r}_{v,y}\in\mathcal{W}^{r}_{new}$ for each edge $e^{r}_{v,y}\in\mathcal{E}^{r}_{new}$ based on the keyword’s importance with in the text corpus and relatedness to the label $y$. We then can merge all graphs $\mathcal{G}^{r}_{new}(\mathcal{V}^{r}_{new},\mathcal{E}^{r}_{new},\mathcal{W}^{r}_{new})$ at each each round $r$ as the full graph $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r},\mathcal{W}^{r})$. More details can refer to Section 3.2.
Part 2: Graph Retrieval The Graph Retrieval process maps the extracted keyword nodes $\mathcal{V}^{t}$ from the unlabeled text $t$ to the constructed information graph $\mathcal{G}^{r}$ and generates a minimum-cost spanning tree on $\mathcal{G}^{r}$ that includes all these keywords. From this minimum-cost spanning tree, the candidate label set $\hat{\mathcal{Y}}_{t}^{r}$ is obtained, which is a reduced subset of the original target label set $\mathcal{Y}^{r}$. For more details, please refer to Section 3.3.
Part 3: Classification After retrieving the graph and obtaining the candidate types $\hat{\mathcal{Y}}_{t}^{r}$ for the text $t$ at round $r$, the classification procedure involves performing the final classification based on a large language model. This process utilizes the text $t$, the candidate labels $\hat{\mathcal{Y}}_{t}^{r}$ retrieved from the information graph $\mathcal{G}^{r}$, and descriptions $\mathcal{K}_{y_{i}}$ associated with each candidate label $y_{i}\in\hat{\mathcal{Y}}_{t}^{r}$. For more details, please refer to Section 3.4.1.
Part 4: Online Indexing The online indexing procedure dynamically indexes keywords, denoted as $\mathcal{V}_{\text{notexist}}^{t}$, which extracted from the testing text $t$ but are not in the existing graph $\mathcal{G}^{r}$. We then integrate these keywords into the graph to further enrich its structure, with weights assigned based on their importance within the text corpus and their relatedness to the predicted label $y^{*}_{t}$ for the testing text $t$. This process enhances the model’s ability to make more accurate predictions in the future. For more details, please refer to Section 3.4.2.
3.2. Part 1: Graph Construction
In this subsection, we would introduce the graph construction procedure of GORAG. To generate the information graph $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r},\mathcal{W}^{r})$ for each the $r$-th round, GORAG applies multiple instructions to the LLM for creating the graph node set $\mathcal{V}^{r}$ and the edges $e^{r}_{v,y}\in\mathcal{E}^{r}$ between keyword node $v$ and label node $y$. Also, we will assign weight $w^{r}_{v,y}\in\mathcal{W}^{r}$ for each each edge $e(v,y)$. The pseudo code of the graph construction algorithm is shown in Algorithm 1.
Algorithm 1 Graph Construction Algorithm of GORAG at the $r$-th round
0:  $\mathcal{D}_{new}^{r}$: Training text set at the $r$-th round and the provided label $y_{i}\in\mathcal{Y}_{new}^{r}$ for each training text $t_{j}$; $p_{extraction},\ p_{gen}$: The extraction and description generation instruction prompt; $\mathcal{G}^{r-1}$: Information graph of the previous round $r-1$.
0:  $\mathcal{G}^{r}$: Constructed information graph at the $r$-th round.
1:  Let $\mathcal{V}_{text}^{r}=\emptyset$, $\mathcal{V}_{label}^{r}=\emptyset$, $\mathcal{E}_{new}^{r}=\emptyset$
2:  for each text $t\in\mathcal{T}_{new}^{r}$ do
3:     $\mathcal{V}_{text}^{r}=\mathcal{V}_{text}^{r}\cup LLM(p_{extract},t)\cup y_{t}$
4:  end for
5:  if label names are available then
6:     for each label $y_{i}\in\mathcal{Y}_{new}^{r}$ do
7:        Get label description $\mathcal{K}_{y_{i}}=LLM(p_{gen},\mathcal{D}^{r})$.
8:        $\mathcal{V}_{label}^{r}=LLM(p_{extract},\mathcal{K}_{y_{i}})\cup\mathcal{V}_{label}^{r}$
9:     end for
10:  end if
11:  Let $\mathcal{V}_{new}^{r}=\mathcal{V}_{text}^{r}\cup\mathcal{Y}^{r}_{new}\cup\mathcal{V}_{label}^{r}$ be the new node set at the $r$-th round.
12:  for each node $v\in\mathcal{V}_{new}^{r}$ do
13:     $\mathcal{E}_{new}^{r}=\mathcal{E}_{new}^{r}\cup e_{v,y_{i}}$
14:     $\mathcal{W}_{new}^{r}=\mathcal{W}_{new}^{r}\cup w_{v,y_{i}}^{r}$
15:  end for
16:  Let $\mathcal{G}^{r}_{new}=(\mathcal{V}_{new}^{r},\mathcal{E}_{new}^{r})$ be the newly constructed graph for at the $r$-th round.
17:  Let $\mathcal{G}^{r}=Merge(\mathcal{G}^{r}_{new},\mathcal{G}^{r-1})$ be the final constructed graph at the $r$-th round.
18:  return  $\mathcal{G}^{r}$
At each $r$-th round, given the labeled training text $\mathcal{D}_{new}^{r}=\cup_{y_{i}\in\mathcal{Y}_{\text{new}}^{r}}\{t_{j},y_{i}\}_{j=1}^{k}$, GORAG first extracts text keywords $\mathcal{K}^{r}_{new}$ from the text $\mathcal{T}^{r}_{new}=\cup_{y_{i}\in\mathcal{Y}_{\text{new}}^{r}}\{t_{j}\}_{j=1}^{k}$ to serve as graph nodes. Specifically, we use the LLM model with an extraction instruction prompt $p_{extract}$, such as $p_{extract}=$Please extract some keywords from the following passage”. Then, we can get all text keywords $\mathcal{K}^{r}_{new}$ based on texts $\mathcal{T}^{r}_{new}$ as follows.
(2) $\mathcal{K}^{r}_{new}=\bigcup_{t\in\mathcal{T}^{r}_{new}}LLM(p_{extract},t).$
Also, by incorporating the new candidate labels $\mathcal{Y}^{r}_{new}$ at the $r$-th round, the graph node set $\mathcal{V}^{r}_{new}$ at the $r$-th round can be obtained as follows.
(3) $\mathcal{V}_{new}^{r}=\mathcal{K}_{new}^{r}\cup\mathcal{Y}_{new}^{r}.$
Then, GORAG links each keyword node $v\in\mathcal{K}_{{new}}^{r}$ to its corresponding label node $y\in\mathcal{Y}_{{new}}^{r}$, indicating that the keyword $v$ appears in texts associated with the label $y$.
We discuss how to compute the weight between each keyword node $v\in\mathcal{K}_{{new}}^{r}$ and each label $y\in\mathcal{Y}_{{new}}^{r}$. Considering the keywords $\mathcal{K}_{new}^{r}$ from the labeled text are not uniformly related to the text’s label $\mathcal{Y}_{new}^{r}$, we apply an edge weighting mechanism to assign a weight $w_{v,y}$ to each keyword-label link. This weight can be regarded as the correlation between keyword and each label. Firstly, we apply the $l_{2}$ normalized TF-IDF score (Sammut and Webb, 2010) to measure the importance and relatedness of a particular keyword $v\in\mathcal{K}_{new}^{r}$ and label $y\in\mathcal{Y}_{new}^{r}$ of the text $t\in\mathcal{T}^{r}$, where $\mathcal{T}^{r}=\cup_{i=1}^{r}\mathcal{T}_{new}^{i}$ denotes texts seen so far. Formally, given a keyword node $v\in\mathcal{K}_{new}^{r}$ and one text $t\in\mathcal{T}^{r}$, the correlation score $CS(v,t)$ between $v$ and $t$ is defined as follows.
(4) $CS(v,t)=\frac{count(v,t)}{|t|}\times log\frac{|\mathcal{T}^{r}|}{1+|t_{j}:v\in t_{j},\ t_{j}\in\mathcal{T}^{r}|},$
where $count(v,t)$ is the number of times that the term $v$ appears in the text $t$, and $|t_{j}:v\in t_{j},\ t_{j}\in\mathcal{T}^{r}|$ denotes the number of texts in the corpus $\mathcal{T}^{r}$ that contain the keyword $v$.
As the keyword $v$ can be extracted from multiple text source and from different rounds, the final edge weight $w_{v,y}^{r}$ of edge $e_{v,y}$ at the $r$-th round is calculated an the average of weights from all these texts. Formally, given the keyword $v\in\mathcal{K}_{new}^{r}$, label $y\in\mathcal{Y}_{new}^{r}$, and the texts $\mathcal{T}^{r,y}=\{t_{i}\}_{i=1}^{|\mathcal{T}^{r,y}|}$ consisting of texts $t_{i}$ associated with label $y$, the weight $w_{v,y}^{r}$ between $v$ and $y$ at the $r$-th round can be computed as follows:
(5) $w_{v,y}^{r}=\frac{\sum_{t_{j}\in\mathcal{T}^{r,y,v}}1-CS(v,t_{j})}{|\mathcal{T}^{r,y,v}|},$
where $\mathcal{T}^{r,y,v}=\{t_{j}|v\in t_{j}\land t_{j}\in\mathcal{T}^{r,y}\}$ is the text associated with label $y$ and containing the keyword $v$. We denote the generated graph for $\mathcal{K}_{new}^{r}$ and $\mathcal{Y}_{new}^{r}$ at the $r$-th round as $\mathcal{G}^{r}_{new}(\mathcal{V}_{new}^{r},\mathcal{E}_{new}^{r},\mathcal{W}_{new}^{r})$, where the node $\mathcal{V}^{r}_{new}=\mathcal{K}_{new}^{r}\cup\mathcal{Y}_{new}^{r}$ and edge $e^{r}_{v,y}\in\mathcal{E}_{new}^{r}$ is associated with the weight $w^{r}_{v,y}\in\mathcal{W}_{new}^{r}$.
We then discuss how to merge $\mathcal{G}^{r}_{new}$ into the graph from previous rounds $\mathcal{G}^{r-1}(\mathcal{V}^{r-1},\mathcal{E}^{r-1})$ to form the full graph $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r})$ as follows.
(6) $\mathcal{G}^{r}=Merge(\mathcal{G}^{r}_{new},\mathcal{G}^{r-1}),$
where $\mathcal{V}^{r}=\mathcal{V}^{r}_{new}\cup\mathcal{V}^{r-1}$, $\mathcal{E}^{r}=\mathcal{E}^{r}_{new}\cup\mathcal{E}^{r-1}$, and $\mathcal{W}^{r}=\mathcal{W}^{r}_{new}\cup\mathcal{W}^{r-1}$. Particularly, $\mathcal{G}^{0}=\emptyset$ and $\ r\geq 1$, Also, to guarantee the graph connectivity of the resulting graph, we add new edges between every newly added label node $y_{n}\in\mathcal{Y}_{new}^{r}$ and each old label node $y_{o}\in\mathcal{Y}^{r-1}$, the edge weight $w_{v,y}^{r}$ of edge $e_{y_{n},y_{o}}$ at round $r$ is calculated with the average weight of all edges that link keywords with label node $y_{n}$ or $y_{o}$, respectively:
(7) $w_{y_{n},y_{o}}^{r}=\frac{AvgWei(y_{n},\mathcal{G}^{r})+AvgWei(y_{o},\mathcal{G}^{r-1})}{2}$
(8) $AvgWei(y,\mathcal{G}^{r})=\frac{\sum_{v\in\mathcal{M}_{y}^{r}}w_{v,y}^{r}}{2\times|\mathcal{M}_{y}^{r}|},$
where $\mathcal{N}^{r}(y)$ denote the neighbor node set of label node $y$ in graph $\mathcal{G}^{r}$. Also, $\mathcal{M}_{y}^{r}=\{v\ |\ v\in\mathcal{N}^{r}(y)\ \land\ v\notin\mathcal{Y}^{r}\}$ denotes all neighbor nodes of label node $y$ that represent keywords, but not labels at the $r$-th round. After merge, the graph $\mathcal{G}^{r}$ would be used for future retrieval, and be further updated by GORAG’s online indexing mechanism later in Section 3.3.
Algorithm 2 Adaptive Candidate Type Generation Algorithm
0:  $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r},\mathcal{W}^{r})$: The constructed information graph; $\mathcal{V}_{exist}^{t}\subseteq\mathcal{V}^{r}$: A set of keyword nodes extracted from text $t$ and can be mapped to graph $\mathcal{G}$; $\mathcal{Y}^{r}$: The target label set at the $r$-th round.
0:  $\hat{\mathcal{Y}}^{r}_{t}$: The candidate type retrieved for text $t$.
1:  Compute a minimum spanning tree $MST$ of the graph $\mathcal{G}$.
2:  Let $\mathcal{V}_{term}=\mathcal{V}_{exist}^{t}\cap\mathcal{V}(MST)$ be the terminal nodes in $MST$
3:  Construct a weighted auxiliary graph $H$: $\mathcal{V}(H)=\mathcal{V}_{term}$
4:  for each pair of terminals $u,v\in\mathcal{V}_{term}$ do
5:     Find the shortest path $P_{uv}$ in $MST$ from $u$ to $v$
6:     Let $w^{H}_{u,v}=\sum_{e\in P_{uv}}w_{e}$
7:  end for
8:  Compute a minimum spanning tree $MST^{\prime}$ of the auxiliary graph $H$
9:  Let $ST=\emptyset$
10:  for each edge $e_{u,v}\in MST^{\prime}$ do
11:     Add the shortest path $P_{uv}$ in $MST$ to $ST$
12:  end for
13:  Let $\hat{\mathcal{Y}}^{r}_{t}=ST\cap\mathcal{Y}^{r}$
14:  return  $\hat{\mathcal{Y}}^{r}_{t}$
3.3. Part 2: Graph Retrieval
In this subsection, we would introduce the graph retrieval procedure of GORAG. With the graph constructed in Part 1, GORAG is able to carry out adaptive reteieval algorithm with the keywords extracted from the testing texts.
To begin with, GORAG extract keywords $\mathcal{V}_{test}^{t}$ for each testing text $t\in T_{test}$ in the same manner with Equation (2), then, $\mathcal{V}_{test}^{t}$ would be splitted into two subsets:
(9) $\mathcal{V}_{exist}^{t}\cup\mathcal{V}_{notexist}^{t}=\mathcal{V}_{test}^{t},\ \mathcal{V}_{exist}^{t}\cap\mathcal{V}_{notexist}^{t}=\emptyset,$
where $\mathcal{V}_{exist}^{t}$ and $\mathcal{V}_{notexist}^{t}$ denotes the keywords in $\mathcal{V}_{test}^{t}$ that already exist and not yet exist in $\mathcal{G}^{r}$ at the current the $r$-th round respectively. Later, $\mathcal{V}_{exist}^{t}$ would be applied for achieving adaptive retrieval, and $\mathcal{V}_{notexist}^{t}$ would be applied for online indexing to further enrich the constructed graph (Further illustrated in Section 3.4).
To achieve the adaptive retrieval, GORAG try to find the minimum cost spanning tree that contain all keyword nodes within $\mathcal{V}_{exist}^{t}$. The intuition behind this approach is that a minimum cost spanning tree (MST) spans the entire graph to cover all given nodes with the smallest possible spanning cost. Consequently, nodes within the generated MST can be considered important for demonstrating the features of the given node set. The generation of an MST is a classical combinatorial optimization problem with an optimal solution determined solely by the set of given nodes. By generating the MST, we eliminate the need for any human-defined thresholds, relying instead on the LLM-extracted keywords to form the spanning tree.
Definition 3.1 (Adaptive Candidate Type Generation Problem).
Given an undirected weighted information graph $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r},\mathcal{W}^{r})$, a set of keywords $\mathcal{V}_{exist}^{t}$ extracted from text $t$ that can be mapped to nodes in $\mathcal{V}^{r}$, and the target label set $\mathcal{Y}^{r}$ at the $r$-th round, our target is to find a set of labels nodes ${\mathcal{Y}}^{r}_{t}\in{\mathcal{Y}}^{r}$. Firstly, we identify a subgraph $\mathcal{G}^{r}_{t}(\mathcal{V}^{r}_{t},\mathcal{E}^{r}_{t},\mathcal{W}^{r}_{t})$ of $\mathcal{G}^{r}$ by minimizing the edge weight sum as follows.
(10) $\min_{e^{r}_{u.v}\in\mathcal{E}^{r}_{t}}w^{r}_{u,v}$
(11) $s.t.$ $v\in\mathcal{V}^{r}_{t},\forall v\in\mathcal{V}_{exist}^{t}$
Then, since the subgraph nodes $\mathcal{V}^{r}_{t}$ contains both keyword nodes and labels, we take the label nodes ${\mathcal{Y}}^{r}_{t}\in\mathcal{V}^{r}_{t}$ as our target candidate nodes for the text $t$.
Theorem 3.2.
The Adaptive Candidate Type Generation problem is NP-hard.
Proof.
To demonstrate that the Adaptive Candidate Type Generation problem is NP-hard, we provide a simple reduction of our problem from the Steiner Tree problem. Since the Steiner Tree problem is already proven to be NP-hard (Su et al., 2020), we show that there is a solution for the Steiner Tree problem if and only if there is a solution for our problem. Firstly, given a solution $S$ for our problem, we can construct a Steiner Tree by generating a minimum spanning tree for all nodes in $S$ on graph $\mathcal{G}^{r}$ then connecting all nodes from $\mathcal{V}_{exist}^{t}$ to their closest neighbor nodes in $S$. Secondly, any Steiner Tree $ST$ that contains any node $y\in\mathcal{V}^{r}_{t}$ is also a solution of our problem with |$|\{y\in\mathcal{V}^{r}_{t}\}|$ nodes. Thus, we prove that the adaptive candidate type generation problem is NP-hard. ∎
Since this problem is the NP-hard, it is infeasible to obtain the optimal result in polynomial time. Therefore, to solve this problem, we propose a greedy algorithm, which generates the candidate type set ${\mathcal{Y}}^{r}_{t}\subseteq\mathcal{Y}^{r}$ for text $t$ at the $r$-th round. The detail of our algorithm is shown in Algorithm 2. Firstly, we calculate the shortest path between each keyword node $v$ in $\mathcal{V}_{exist}^{t}$ to all other nodes in $\mathcal{G}^{r}$ by calculating the minimum spanning tree $MST$ of $\mathcal{G}^{r}$ (line 1-2). Here, the keyword nodes $v$ in $\mathcal{V}_{exist}^{t}$ are served as terminal nodes that determins the final genrated candidate types w.r.t. to the information graph $\mathcal{G}^{r}$ Secondly, we create a new auxiliary graph $H$ where the edges represent the shortest paths between the closest terminal nodes (line 3-7). Thirdly, we construct the minimum spanning tree $MST^{\prime}$ of the auxiliary graph $H$ (line 9), and then we add the shortest paths between each two nodes in $MST^{\prime}$ to the Steiner Tree $ST$ (line 10-12). Lastly our candidate types $\hat{\mathcal{Y}}^{r}_{t}$ for text $t$ are calculated as the interselect of $ST$ and all target labels $\mathcal{Y}^{r}$ at round $r$: $ST\cap\mathcal{Y}^{r}$ (line 13).
Note that in GORAG’s retrieval algorithm, the minimum-cost spanning tree is generated solely based on the constructed graph and the keywords extracted from the text. This approach eliminates the need for any manually defined retrieval thresholds, enabling GORAG to perform fully adaptive and context-driven retrieval.
3.4. Part 3: Classification and Online Indexing
In this subsection, we introduce GORAG’s two main components, i.e., classification and online indexing. For classification, GORAG utilizes an LLM to predict the class label for each unseen test text by retrieving adaptive demonstrations. For online indexing, GORAG dynamically updates the graph $\mathcal{G}^{r}$ by incorporating unseen keywords from the test text $t$ as new nodes and connecting them to the predicted label $y^{*}_{t}$ with weighted edges. This iterative process enriches the graph and enhances the framework’s ability to make accurate predictions in the future.
3.4.1. Classification
In this part, we introduce how GORAG performs text classification based a large language model. Specifically, for each unlabeled text $t$, GORAG predicts its class label $y^{*}$ by constructing an input $c_{t}$ for the LLM, which captures contextual and structural information about the text and its candidate labels. Specifically, given each unlableed text $t$, we will use LLM to predict its class label $y^{*}$. We first generate the LLM input $c_{t}$ as follows.
(12) $c_{t}=Concat(t,\mathcal{K}^{t},\hat{\mathcal{Y}}^{r}_{t},\mathcal{K}_{\hat{\mathcal{Y}}^{r}_{t}}),$
which is the text concatenation of the extracted keywords $\mathcal{K}^{t}$ from the text $t$, the candidate labels $\hat{\mathcal{Y}}^{r}_{t}$ obtained by Algorithm 2. Also, $\mathcal{K}_{\hat{\mathcal{Y}}^{r}_{t}}=\{\mathcal{K}_{y_{i}}\}_{i=1}^{|\hat{\mathcal{Y}}^{r}_{t}|}$ and $\mathcal{K}_{y_{i}}$ is the representative keywords of each label $y_{i}\in\hat{\mathcal{Y}}^{r}_{t}$.
(13) $\mathcal{K}_{y_{i}}=LLM(p_{gen},\mathcal{D}^{r}),$
and a classification instruction prompt $p_{classify}$.
(14) $y^{*}_{t}=LLM(p_{classify},c_{t}),$
As a result, the LLM would try to select the best-suited label $y^{*}_{t}\in\mathcal{Y}^{t}$ to annotate the text $t$.
3.4.2. Online indexing
To fully leverage the text-extracted keywords, GORAG utilizes an online indexing mechanism to incrementally update keywords that do not yet exist in the information graph $\mathcal{G}^{r}$ at the $r$-th round to the information graph $\mathcal{G}^{r}$ based on the text $t$’s predicted label $y^{*}_{t}$. To be specific, each keyword node $v\in\mathcal{V}_{notexist}^{t}$ would be directed added to the original graph’s node set $\mathcal{V}^{r}$ and be assigned with an edge $e_{v,y}$ connect it with the predicted label $y$:
(15) $\mathcal{V}^{r}=\mathcal{V}^{r}\cup\mathcal{V}_{notexist}^{t},\ \mathcal{E}^{r}=\mathcal{E}^{r}\cup\mathcal{E}_{oi}^{t},$
where $\mathcal{E}_{oi}^{t}=\{e_{v,y_{t}}|v\in\mathcal{V}_{notexist}^{t}\}$ denotes the set of all newly assigned edges $e_{v,y}$ between keyword node $v$ and its predicted label $y_{t}^{*}$, for these newly assigned edges, their weight is calculated with the edge weighting mechanism illustrated in Equation (5).
Table 2. Statistics of the WOS and Reuters dataset.
Dataset Avg. Text Token R1 R2 R3 R4 Total
Testing data Label # Testing data Label # Testing data Label # Testing data Label # Testing data
WOS (Kowsari et al., 2017) 200 2,417 32 2,842 53 2,251 30 1,886 18 9,396
Reuters (Bao et al., 2020) 168 80 8 80 8 80 8 70 7 310
This online indexing mechanism serves two main purposes. First, it incrementally enriches the information graph $\mathcal{G}^{r}$ by incorporating new, previously unseen keywords, thus expanding the graph’s vocabulary and improving its adaptability. Second, by linking these new keywords to their predicted labels with weighted edges, the mechanism captures their relevance and context more effectively, enabling better utilization of the graph for future predictions.
3.5. Complexity Analysis
In this subsection, we analysis the time and space complexity of GORAG’s graph construction, retrieval and prediction procedure. We denote the maximum number of terms of the input text as $m_{t}$, the number of unique terms of training corpus $\mathcal{D}_{new}^{r}$ as $u$, the LLM’s maximum input token length as $m_{i}$, LLM’s its maximum extraction, generation, and classification token length as $m_{e}$, $m_{g}$, and $m_{c}$, respectively.
Graph Construction Complexity Firstly, we analysis the time and space complexity of GORAG’s graph construction. For the text keyword extraction procedure, the time complexity would be $O(|\mathcal{D}_{new}^{r}|(m_{i}+m_{e}))$; If label names are available, the time complexity of generat label descriptions and extract label keywords would be $O(|\mathcal{Y}^{r}_{new}|(m_{i}+m_{g}+m_{e}))$; Calculating the TFIDF and indexing edges to graph would require $O(|\mathcal{V}_{new}^{r}|(|\mathcal{D}_{new}^{r}|m_{t}+u))$ times, and merging graph requires $O(|\mathcal{Y}^{r}_{new}||\mathcal{Y}^{r-1}|)$. Hence, the total time complexity of GORAG’s graph construction at the $r$-th round would be:
$O(|\mathcal{D}_{new}^{r}|(m_{i}+m_{e})+|\mathcal{Y}^{r}_{new}|(m_{i}+m_{g}+m_{e}+|\mathcal{Y}^{r-1}|)+|\mathcal{V}_{new}^{r}|(|\mathcal{D}_{new}^{r}|m_{t}+u))$.
For the space complexity, the graph is stored with the weighted adjency matrix, hence needs $O(|\mathcal{E}^{r}|)$ space; Storing the training corpus at the $r$-th round would need $O(u|\mathcal{D}_{new}^{r}|)$ space; Storing the representive keywords $\mathcal{K}_{\hat{\mathcal{Y}_{t}^{r}}}$ would need $O(|\mathcal{K}_{\hat{\mathcal{Y}_{t}^{r}}}|)$ space. Hence the total space complexity of GORAG’s graph construction at the $r$-th round would be:
$O(|\mathcal{E}^{r}|+u|\mathcal{D}_{new}^{r}|+|\mathcal{K}_{\hat{\mathcal{Y}_{t}^{r}}}|)$.
Retrieval and Classification Complexity The time complexity of GORAG’s adaptive candidate type generation algorithm is the same with the Mehlhorn algorithm, which is $O(|\mathcal{E}^{r}|+|\mathcal{V}^{r}|log|\mathcal{V}^{r}|)$ (Mehlhorn, 1988); The time complexity of the online indexing mechanism would cost $|\mathcal{V}_{notexist}^{r}|(|\mathcal{T}_{test}^{r}|m_{t}+u)$; The time complexity of the final classification by LLM is $O(m_{i}+m_{c})$. Hence, the total time complexity of GORAG’s adaptive retrieval and classification can be denoted as
$O(|\mathcal{T}_{test}^{r}|(|\mathcal{E}^{r}|+|\mathcal{V}_{notexist}^{r}|(|\mathcal{T}_{test}^{r}|m_{t}+u)+|\mathcal{V}^{r}|log|\mathcal{V}^{r}|+m_{i}+m_{c})).$
For the space complexity, $O(|\mathcal{E}^{r}|)$ to store the graph, and $O(u|\mathcal{T}_{test}^{r}|)$ space is needed to store the testing corpus. Hence, the total space complexity of GORAG’s Retrieval and Classification procedure is
$O(|\mathcal{E}^{r}|+u|\mathcal{T}_{test}^{r}|).$
4. Experiments
In this section, we present the experimental evaluation of our framework GORAG on the DFSTC task. We compare the performance of GORAG against six effective baselines spanning three technical categories, using two datasets with distinct characteristics. Specifically, we first outline the experimental setup in Section 4.1, including details on datasets, baselines, evaluation metrics, and hyperparameter configurations. Next, we report the experimental results, focusing on both effectiveness and efficiency evaluations, in Section 4.2. Finally, we conduct an ablation study and a case study, presented in Section 4.3 and Section 4.4, respectively.
Table 3. Experiment results on WOS and Reuters dataset.
Dataset Category Model Round
R1 R2 R3 R4
1-shot 5-shot 10-shot 1-shot 5-shot 10-shot 1-shot 5-shot 10-shot 1-shot 5-shot 10-shot
WOS NN-based Entailment 0.3695 0.3823 0.4187 0.3994 0.4471 0.4222 0.4510 0.4857 0.4787 0.4030 0.4387 0.4442
Long Context RAG NaiveRAG 0.3885 0.3904 0.3897 0.2267 0.2154 0.2187 0.1821 0.1475 0.1799 0.1653 0.1556 0.1649
Propositionizer 0.1241 0.1306 0.1297 0.1074 0.1421 0.1303 0.1771 0.1645 0.1603 0.1611 0.1637 0.1656
Compression RAG LongLLMLingua 0.3806 0.3823 0.3901 0.2155 0.2202 0.2198 0.1770 0.1567 0.1608 0.1468 0.1382 0.1493
GraphRAG 0.3852 0.3897 0.3906 0.2213 0.2197 0.2219 0.1816 0.1770 0.1786 0.1641 0.1634 0.1625
LightRAG 0.3930 0.3806 0.3815 0.2202 0.2216 0.2145 0.1743 0.1767 0.1799 0.1625 0.1626 0.1632
GORAG 0.4862 0.4973 0.5192 0.4649 0.5063 0.5208 0.4814 0.4906 0.4993 0.4210 0.4340 0.4471
Reuters NN-based Entailment 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
Long Context RAG NaiveRAG 0.0000 0.0000 0.0375 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
Propositionizer 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
Compression-RAG LongLLMLingua 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0258 0.0065 0.0085
GraphRAG 0.1375 0.1625 0.1000 0.0688 0.0813 0.0500 0.0375 0.0417 0.0417 0.0291 0.0375 0.0375
LightRAG 0.0500 0.1375 0.1125 0.0250 0.0813 0.0563 0.0333 0.0333 0.0333 0.0125 0.0333 0.0208
GORAG 0.0875 0.0875 0.1000 0.1750 0.1688 0.1438 0.1667 0.1958 0.2167 0.1667 0.2516 0.2193
4.1. Experiment Settings
4.1.1. Datasets
We select the Web-Of-Science (WOS) dataset (Kowsari et al., 2017) and the Reuters few-shot text classification dataset by (Bao et al., 2020) for evaluating GORAG’s performance. We split the WOS dataset into 4, 6 and 8 rounds, and the Reuters dataset into 4 rounds, while trying to maintain a comparable number of testing data in each round. Due to the spacec limitation, we mainly display our experinents on the 4 rounds spit version of both datasets.
For the WOS dataset, the texts to be classified are text chunks from academic research papers, and their labels are the respective research fields. In the original WOS dataset, there are 9,396 testing texts and 133 labels with label names avaiable for all rounds. For Reuters datasets, the texts to be classified are text chunks of news, and their labels are numbers representing the news categories. In the original Reuters text classification dataset, there are 310 testing texts and 31 labels for all rounds, and the label names are not avaiable. Further statistics of these two dataset being splitted into 4 rounds are shown in Table 2.
Considering the differences in the properties of the two datasets, conducting experiments on both datasets provides a comprehensive and in-depth understanding of GORAG’s characteristics.
4.1.2. Baselines
In this paper, we compare GORAG’s performance with 7 baselines from 3 technical categories:
NN-based Models
  • Entailment (Xia et al., 2021): Entailment concatenates the text data with each of the label names to form multiple entailment pairs with one text sample, hence increasing the number of training data and enhance its finetuning of a RoBERTa PLM (Liu, 2019). Entailment then carry out classification based on these entailment pairs and convert the text classification task into a binary classification task, for detecting whether the entailment pair formed is correct.
Long Context RAG Models
  • NaiveRAG (Gao et al., 2023): NaiveRAG acts as a foundational baseline of current RAG models. When indexing, it stores text segments of the labeled texts in a vector database using text embeddings. When querying, NaiveRAG generates query texts’ vectorized representations to retrieve side information based on the highest similarity in their embeddings.
  • Propositionizer (Chen et al., 2023): Propositionizer applies a fine-tuned LLM to convert side information into atomic expressions, namely propositions, to facilitate more fine-grained information retrieval than NaiveRAG.
Compression RAG Models
  • LongLLMLingua (Jiang et al., 2024): LongLLMLingua is a instruction aware prompt compressor model, it applies LLM’s generation perplexity to filter out un-important tokens of the model input based on the retrieved side information and the task instruction.
  • GraphRAG (Edge et al., 2024): GraphRAG is a graph-Based RAG model that employs the LLM to extract texts’ entities and relations, which are then represented as nodes and edges in the information graph. GraphRAG then aggregates nodes into communities, and generates a comprehensive community report to encapsulate global information from texts.
  • LightRAG (Guo et al., 2024): LightRAG skips the GraphRAG’s formulation of graph communities and directly retrieve nodes or pathes from created graphs. It calculates the embedding similarity of text extracted entities with graph nodes, to achieve a one-to-one mapping from keywords to graph nodes and retrieve all triples involved these nodes.
According to Table 3, among these methods, Entailment achieves the state-of-the-art performance on the DFSTC task. For GraphRAG and NaiveRAG, we use the implementation from (nan, [n.d.]) which optimizes their original code and achieves better time efficiency while not affect the performance; For all other baselines, we use their open-sourced official implementations.
4.1.3. Evaluation Metrics
In this paper, we use classification accuracy as the evaluation metric. Given that LLMs can generate arbitrary outputs that may not precisely match the provided labels, we consider a classification correct only if the LLM’s output exactly matches the ground-truth label name or label number.
(a) 1-shot setting.
(b) 5-shot setting.
Figure 4. Experiment result of WOS dataset under 1-shot and 5-shot settings.
(a) Model Training/Indexing time on WOS dataset.
(b) Model Training/Indexing time on Reuters dataset.
Figure 5. Model Training/Indexing time on two datasets under 1-shot setting.
4.1.4. Hyperparameter and Hardware Settings
In each round of our experiments, we handle NN and RAG-based models differently for training and indexing. For NN-based models, we train them using all labeled data from the current round and all previous rounds. For RAG-based models, we index the labeled data of the current round to the information source from the previous round. The RAG-based model is initialized from scratch only in the first round.
After training or indexing in each round, we test the models using the testing data from the current round and all previous rounds. For GORAG, which employs an online indexing mechanism, we first test its performance with online indexing on the current round’s testing data. We then test its performance on the testing data from all previous rounds without online indexing. This approach allows us to study the effect of online indexing on the performance of later round data when applied to previous round data.
For the hyperparameter settings, we train the Entailment model with the RoBERTa-large PLM for 5 epochs with a batch size of 16, a learning rate of $1\times 10^{-6}$, and the Adam optimizer (Kingma, 2014), following the exact setting in Entailment’s original paper. For LongLLMLingua, we test the compression rate within ${0.75,0.8,0.85}$ and select 0.8, as it achieves the best overall classification accuracy. For GraphRAG and LightRAG, we use their local search mode, as it achieves the highest classification accuracy on both the WOS and Reuters datasets. For GORAG and all RAG-based baselines, we use LLaMA3 (Touvron et al., 2023) as the LLM backbone unless otherwise specified.
All experiments are conducted on an Intel(R) Xeon(R) Gold 5220R @ 2.20GHz CPU and a single NVIDIA A100-SXM4-40GB GPU.
4.2. Experiment Results
In this paper, we employ 1-shot, 5-shot, and 10-shot settings for few-shot training, where each setting corresponds to using 1, 5, and 10 labeled training samples per class, respectively. We omit experiments with more than 10 labeled samples per class, as for WOS dataset, there are already over 1300 labeled training data under 10-shot setting.
4.2.1. Effectiveness Evaluation
As shown in Table 3, GORAG achieves the best classification accuracy over all four rounds with 1-shot and 5-shot labeled indexing data on WOS dataset, it surpasses all RAG-based baselines as well as the state-of-the-art model Entailment. Compared with Entailment, when apply 1-shot setting, GORAG achieves at most 31.6% accuracy gain from 0.3695 to 0.4862 at the first round. On Reuters dataset, GORAG achieves the best classification accuracy over the last 3 rounds with 1-shot indexing data.
Furthermore, based on the experiment results in Table 3, we can have following observations. Firstly, for RAG-based models other than GORAG, although they can achieve comparable or even better performance than NN-based models in the first round, they tend to suffer a more significant performance drop as the number of labels increases. This is because the current RAG models compress the texts and ignore the compression of the target label set. As a result, when the number of target labels increases, the lengthy input can make the LLM tend to give wrong classification results and suffer from the lost-in-the-middle issue.
Secondly, compared with NN-based models, all RAG-based models other than GORAG are less sensitive to the increase of labeled data, from 1-shot to 5-shot setting on WOS dataset, the performance of Entailment increases significantly, however, for RAG-based baselines, performance tends to remain the same from 1-shot to 10-shot setting. This is because RAG-based models do not train any model parameters. Although an increase in training data can bring new information to enrich the side information, they are not always considered useful by the model and may not be retrieved. Consequently, the increase in labeled data may have no effect on some data samples during testing. On the other hand, for GORAG, the new information can introduce new paths for GORAG’s adaptive retrieval algorithm. This expands the influence scope of the new information from a fixed graph neighborhood to the entire information graph, making it easier for the retrieval process to benefit from the new information.
Thirdly, for Long Context RAG models, their lengthy retrieved side information further increase the length of LLM classification inputs, and consequently make the LLM harder to detect the relevence between side information and the labels than Compression based RAG models, exacerbating the lost-in-the-middle issue and make the LLM tend to generate wrong classification results.
4.2.2. Robustness Evaluation
To evaluate the robustness of GORAG and several representative baselines from the three technical categories, we conducted additional experiments on the 6 and 8 round split versions of the WOS dataset under 1-shot and 5-shot settings, as the WOS dataset contains more testing data than Reuters dataset. In Figure 4, the results for rounds 1 to 4 are obtained from the 4-round split version, rounds 5 and 6 are obtained from the 6-round split version, and rounds 7 and 8 are obtained from the 8-round split version of the WOS dataset. This setting ensures that the testing data at each round is as sufficient as possible while maintaining a acceptable amount of experiment time cost. As illustrated in Figure 4, the classification accuracy of GraphRAG and NaiveRAG drops significantly from round 1 to round 8, demonstrating the negative impact of the lengthy and unfiltered target label set on classification results. However, compared to other RAG-based baselines, GORAG maintains competitive classification accuracy as the number of rounds increases through 1 to 8 round, with both of the 1-shot and 5-shot setting.
Table 4. The size of the constructed information graph in each round.
Dataset Model Round
R1 R2 R3 R4 After R4
Node # Edge # Node # Edge # Node # Edge # Node # Edge # Node # Edge #
WOS GORAG offline 1,149 1,405 3,681 4,587 4,947 6,204 5,438 6,879 5,438 6,879
GORAG 1,149 1,405 12,108 12,245 25,357 26,290 34,521 35,919 44,283 45,973
Reuters GORAG offline 117 119 192 200 263 284 331 366 331 366
GORAG 117 119 584 594 1,013 1,046 1,342 1,405 1,533 1,600
Table 5. 0-shot experiment result on WOS dataset.
Model Round
R1 R2 R3 R4
Qwen2.0-7B 0.3322 0.1790 0.1451 0.1444
Mistral0.3-7B 0.1436 0.0540 0.0336 0.0270
LLaMA3-8B 0.3351 0.1614 0.1161 0.0930
GORAG 0.3305 0.2567 0.2230 0.2102
4.2.3. Efficiency Evaluation
To evaluate the efficiency of GORAG, we compare GORAG’s indexing time cost with other baselines’ training or indexing time cost on the 4 round split version of WOS and Reuters datasets, the results are shown in Figure 5. For both datasets, Propositionizer achieves the worst indexing efficiency. This is because it employs an additional fine-tuned LLM to convert side information into propositions. In this process, this extra LLM generates a long list of propositions converted from each indexing or testing text, introducing more efficiency overhead compared to other models that use the LLM solely for classification, which typically generate only a few tokens representing the label name or label number.
On the WOS dataset, the indexing procedure of compression RAG-based models exhibits better time efficiency than the training procedure of NN-based models from R2 to R4. This is because NN-based models require retraining the entire model with all seen labeled data in each round, which limits their training efficiency.
On the Reuters dataset, GraphRAG achieves the worst indexing efficiency among all Compression RAG models. This is because the Reuters dataset contains fewer labels than the WOS dataset in each round. As a result, the constructed graph tends to be smaller than that of the WOS dataset, as shown in Table 4. Consequently, since the time cost of GraphRAG’s community generation procedure is also determined by the number of communities generated (Traag et al., 2019), GraphRAG introduces a more significant efficiency overhead on the Reuters dataset where the created information graph is much smaller compared to the WOS dataset.
4.3. Ablation Study
To further study GORAG’s zero-shot ability, we conducted an experiment on the WOS dataset by providing only the label names, without any labeled data to GORAG and some widely applied open-source LLM models, the result is shown in Table 5. In the 0-shot setting, GORAG first generates label descriptions based on the label names, then extracts keywords from these descriptions, without using any information from labeled texts. Note that the Entailment model cannot be applied in this setting, as they require labeled data for training or indexing. For RAG-based models, their performance in the 0-shot setting is equivalent to that of their backbone LLMs due to the lack of indexing data.
As shown by the results, GORAG achieves comparable performance with other open-source LLM models in the first round and outperforms them in all subsequent rounds. This is because, when the number of target labels is small, the benefit of compressing the target label set is not as pronounced. However, as the number of labels increases, the importance of filtering the label set becomes more significant, allowing GORAG to consistently outperform other open-source LLM models.
To better study the reason behind GORAG’s superior performance, we conduct an extensive ablation study under 1-shot setting on GORAG. For the ablation studie, we only experimented with the 1-shot setting unless further illustrated, as the performance of RAG models do not change significantly with different shot settings. Specifically, to better study how different components of GORAG affects the classification performance, we apply its following variants for ablation study:
  • GORAG unit: The GORAG model that remove the edge weighting mechanim, every edge in this variant is assigned with weight 1.
  • GORAG offline: The GORAG model that remove the online indexing mechanism.
  • GORAG keyword: The GORAG model that only use the keyword extracted from the text to create LLM classification input, rather than the whole text.
Table 6. Ablation experiments for variants of GORAG.
Dataset Model Round
R1 R2 R3 R4
WOS GORAG unit 0.4706 0.4394 0.4407 0.3899
GORAG offline 0.3063 0.2302 0.2455 0.2156
GORAG keyword 0.4746 0.4606 0.4455 0.4030
GORAG 0.4862 0.4649 0.4814 0.4210
Reuters GORAG unit 0.0000 0.0625 0.1667 0.0795
GORAG offline 0.0000 0.1000 0.0917 0.1032
GORAG keyword 0.0000 0.1000 0.1167 0.1438
GORAG 0.0875 0.1750 0.1667 0.1667
(a) GORAG’s candidate selection algorithm returns candidate labels adaptively based on keywords extracted from the tesing texts.
(b) GORAG’s candidate selection algorithm only returns the ground truth label of the testing text.
Figure 6. Two case of GORAG successfully generate candidate labels to help text classification on WOS dataset.
In Table 6, we present the results of these variants to evaluate whether the edge weighting mechanism, the online indexing mechanism, and the substitution of lengthy texts with shorter keywords can enhance the prediction accuracy of GORAG. The results show that GORAG achieves the best prediction accuracy, demonstrating the importance of the edge weighting and onling indexing mechanism. Furthermore, GORAG keyword also achieves a competitive performance, illustrating the keywords extracted from texts can also have valuable information for text classification.
Table 7. Evaluation on different LLM backbones.
Dataset Model Round
R1 R2 R3 R4
WOS GORAG LLaMA3 0.4862 0.4649 0.4814 0.4210
GORAG Qwen2.5 0.5101 0.4839 0.4823 0.4235
Additionally, in Table 4, we compare the size of the constructed information graph with and without online indexing after each round. After applying the online indexing mechanism, the information graph being significantly enriched from R1 to After R4.
In Table 7, we present the results of GORAG with different backbone LLMs, we test GORAG with more advanced Qwen2.5-7B (Qwen Team, 2024), the result demonstates the GORAG’s performance can be further improved with more advanced LLM backbones.
4.4. Case Study
In this subsection, to illustrate the strength of GORAG’s adaptive retrieval and candidate label generation procedure, we dig into two testing cases select from WOS dataset where GORAG’s adaptive retrieval helps to reduce the LLM input length and benefit text classification. As shown in Figure 6(a), for the case whose ground truth label is Algorithm Design, the candidate label retrieved by GORAG are Computer programming and Algorithm design. As a result, GORAG’s adaptive retrieval algorithm successfully filter the target label set to only contains two candidate labels, and successfully cut down the target label number from at most over 100 to only 2 candidate labels that it consided as the possible labels for the text, significantlly reduce the LLM’s input length and mitigate the lost-in-the-middle issue.
On the other hand, for the case in Figure 6(b) whose ground truth label is Electrical Circuits, since the ground truth label name Electrical Circuits already exists in the text and being extracted as keyword, GORAG’s adaptive retrieval algorithm is almost certain to classify this text into class Electrical Circuits, hence GORAG’s adaptive retrieval and candidate type generation algorithm would only select the only candidate label retrieved from the constructed graph and input to the LLM.
4.5. Discussion
According to the aforementioned experimental results, we have identified several key characteristics of NN-based models and RAG-based models.
On the one hand, despite the rapid development of LLMs in recent years, they can still yield sub-optimal classification accuracy in DFSTC tasks due to the lengthy LLM inputs. When there is a sufficient amount of labeled data available in each dynamic updation round, the NN-based model Entailment can also achieve comparable or even better performance than traditional RAG based models other than GORAG.
On the other hand, compared with NN-based models, benefited from the time efficient indexing of the side information and the rich pre-trained knowledge within the LLM, by filter out some unrelated labels and shorten the LLM inputs, RAG-based models are better suited for DFSTC task when the target labels are updated frequently and there only existes a limit number or even none of the labeled data is avaiable for each dynamic updation round.
5. Conclusion
In this paper, we propose GORAG, a Graph-based Online Retrieval Augmented Generation framework for the Dynamic Few-shot Text Classification (DFSTC) task. Extensive experiments on two text datasets with different characteristics demonstrate the effectiveness of GORAG in classifying texts with only a limited number or even no labeled data. Additionally, GORAG shows its effectiveness in adapting to the dynamic updates of target labels by retrieving candidate labels to filter the large target label set in each update round. With extensive of ablation studies, we confirm that GORAG’s adaptive retrieval, edge weighting, and online indexing mechanisms contribute to its effectiveness. For future work, we aim to further enhance GORAG’s performance and explore its application in more complex and diverse scenarios, as well as further improve the efficiency of GORAG’s online indexing mechanism.

### v2

@@
-Graph-based Retrieval Augmented Generation for Dynamic Few-shot Text Classification
+GORAG: Graph-based Retrieval Augmented Generation for Dynamic Few-shot Text Classification
 Abstract.
-Text classification is a fundamental task in natural language processing, pivotal to various applications such as query optimization, data integration, and schema matching.
-While neural network-based models, such as CNN and BERT, have demonstrated remarkable performance in text classification, their effectiveness heavily relies on abundant labeled training data.
-This dependency makes these models less effective in dynamic few-shot text classification, where labeled data is scarce, and target labels frequently evolve based on application needs.
-Recently, large language models (LLMs) have shown promise due to their extensive pretraining and contextual understanding.
-Current approaches provide LLMs with text inputs, candidate labels, and additional side information (e.g., descriptions) to predict text labels.
+Text classification is a fundamental task in data mining, pivotal to various applications such as tabular understanding and recommendation.
+Although neural network-based models, such as CNN and BERT, have demonstrated remarkable performance in text classification, their effectiveness heavily relies on abundant labeled training data.
+This dependency makes these models less effective in dynamic few-shot text classification, where labeled data is scarce, and new target labels frequently appear based on application needs.
+Recently, large language models (LLMs) have shown promise due to their extensive pretraining and contextual understanding ability.
+Current approaches provide LLMs with text inputs, candidate labels, and additional side information (e.g., descriptions) to classify texts.
 However, their effectiveness is hindered by the increased input size and the noise introduced through side information processing.
 To address these limitations, we propose a graph-based online retrieval-augmented generation framework, namely GORAG, for dynamic few-shot text classification.
-GORAG constructs and maintains an adaptive information graph by extracting side information across all target texts, rather than treating each input independently.
-It employs a weighted edge mechanism to prioritize the importance and reliability of extracted information and dynamically retrieves relevant context using a minimum-cost spanning tree tailored for each text input.
-Empirical evaluations demonstrate that GORAG outperforms existing approaches by providing more comprehensive and accurate contextual information.
-PVLDB Reference Format:
-PVLDB, 14(1): XXX-XXX, 2020.
-doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License.
-Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license.
-For any use beyond those covered by this license, obtain permission by emailing info@vldb.org.
-Copyright is held by the owner/author(s).
-Publication rights licensed to the VLDB Endowment.
-Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
-doi:XX.XX/XXX.XX
-PVLDB Artifact Availability:
-The source code, data, and/or other artifacts have been made available at ***.
+Rather than treating each input independently, GORAG constructs and maintains a weighted graph by extracting side information across all target texts.
+In this graph, text keywords and labels are represented as nodes, with edges indicating the correlations between them.
+To model these correlations, GORAG employs an edge weighting mechanism to prioritize the importance and reliability of extracted information and dynamically retrieves relevant context using a minimum-cost spanning tree tailored for each text input.
+Empirical evaluations demonstrate that GORAG outperforms existing approaches by providing more comprehensive and precise contextual information.
+Large language models, online retrieval-augmented generation, few-shot learning
+copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2025; Woodstock, NYisbn: 978-1-4503-XXXX-X/2025/02ccs: Information systems Clustering and classificationccs: Information systems Retrieval models and ranking
 1. Introduction
-Text classification is a fundamental task that is connected to various other tasks, such as query optimization (Yu et al., 2023; Yu and Litchfield, 2020) data integration (Dong and Rekatsinas, 2018), and schema matching (Sahay et al., 2020; Zhang et al., 2024a).
-For example, in schema matching tasks, a text classifier can be trained to determine whether two schema attributes refer to the same concept based on their descriptions or metadata (Zhang et al., 2024a).
+Text classification is a fundamental task of text data mining and is connected to various tasks, such as tabular understanding (Chen et al., 2024c; Wang et al., 2024c; Wen et al., 2024a, b; He et al., 2025; Zheng et al., 2023), and recommendation (Du et al., 2024a; Han et al., 2024b; Wang et al., 2024b; Li et al., 2023; Zheng et al., 2024; Chiang et al., 2023; Du et al., 2024b).
 In recent years, Neural Network (NN)-based models (Johnson and Zhang, 2015; Prieto et al., 2016; Liu et al., 2016; Dieng et al., 2017; Wang et al., 2017; Johnson and Zhang, 2017; Peters et al., 2018; Wang et al., 2018; Yang et al., 2019; Miao et al., 2021), such as CNN (Kim, 2014), Bert (Devlin, 2018) and RoBERTa (Liu, 2019), have demonstrated impressive performance on text classification tasks.
-However, the effectiveness of these NN-based approaches depend on abundant labeled training data, which requires significant time and human efforts (Meng et al., 2018).
-Consequently, these methods perform poorly with limited labeled text data (Xia et al., 2021), such as in few-shot text classification tasks (Wang et al., 2020; Aljehani et al., 2024).
+However, the effectiveness of these NN-based approaches depends on abundant labeled training data, which requires significant time and human efforts (Meng et al., 2018).
+Furthermore, in real-world applications, such as Web Of Science (Kowsari et al., 2017) and IMDb (imd, [n.d.]), target labels for text often change based on the application’s requirements (Xia et al., 2021), leading to the dynamic few-shot text classification (DFSTC) task.
+As shown in Figure 1, the DFSTC task begins with a few initial classes (e.g., Machine Learning).
+As new topics like Computer Vision emerge in later rounds, models must adapt to these changes and accurately classify new examples with minimal labeled data.
+Hence, how to develop methods to dynamically classify text with limited labeled data available remains a necessity and open problem.
+Figure 1. An example of the DFSTC task with two rounds. In round 1, an example text is classified as Machine Learning. In round 2, an example text is classified as Computer Vision.
+Depending the technique for the DFSTC task, current models can be mainly categorized into two types, i.e., data augmentation-based models and RAG-based models.
+Firstly, data augmentation-based models (Meng et al., 2018; Meng et al., 2020; Xia et al., 2021) create additional training data by mixing the pairs of the few-shot labeled data and assigning mixed labels indicating the validity for these created data based on the labels of each data pairs (Xia et al., 2021).
+Also, several data augmentation approaches create extra semantic-related content based on the label names (Meng et al., 2018; Meng et al., 2020) to enrich the training corpus further.
+However, due to the limited labeled data, and the label names may not always be available, the generated text data can have very limited patterns.
+Consequently, the text classification models trained on these generated data may over-fit on limited text data and are not generalizable (Liu et al., 2023; Liu et al., 2024c; Lei et al., 2023).
 Figure 2.
-An illustration of the structural composition of current Reteieval Augmented Generation based few-shot text classification models, where the few-shot labeled training data is indexed as an external information source.
-During classification, the model retrieves side information from the external information source to provide additional context that aids the classification process.
-Also, in real-world applications, such as Web Of Science (Kowsari et al., 2017) and IMDb (imd, [n.d.]), target labels for text often change based on the application’s requirements (Xia et al., 2021), leading to dynamic text classification.
-For instance, movie labels might initially be {Action, Thriller, Comedy}, but platforms may need to add more genres (e.g., Romantic and Comic) to extend the original labels.
-Therefore, it is important to classify text into these new categories.
-Consequently, how to develop methods to dynamically classify text with limited labeled data (i.e., dynamic few-shot text classification) remains a necessity and open problem.
-Depending the technique for dynamic few-shot text classification task, current models can be categorized into three types, i.e., data augmentation-based (Meng et al., 2018; Meng et al., 2020; Xia et al., 2021), meta learning-based (Bansal et al., 2020; Zhang et al., 2023a), and large language model (LLM)-based approaches (Chen et al., 2023; Sarthi et al., 2024; Edge et al., 2024; Gutiérrez et al., 2024; Guo et al., 2024).
-Firstly, data augmentation-based models create additional training data by mixing the pairs of the few-shot labeled data and assign a mixed labels for these created data based on the labels of each pairs of data (Xia et al., 2021).
-Also, several data augmentation approaches create extra semantic related content based on the label names (Meng et al., 2018; Meng et al., 2020).
-However, due to the limited labeled data, the generated text data has very limited patterns.
-Consequently, the text classification models trained on these generated data are over-fitting on limited text data and are not generalizable (Liu et al., 2023).
-Also, meta-learning models (Bansal et al., 2020; Zhang et al., 2023a) train meta-learners on each base class to quickly adapt to new classes with minimal data.
-However, meta-learning still requires a substantial amount of labeled data for base classes.
-Recently, large language models (Sun et al., 2021; Jiang et al., 2023; Touvron et al., 2023; Achiam et al., 2023; Yang et al., 2024), pretrained on extensive corpora, have achieved significant success due to their superior and comprehensive text understanding abilities.
-Several researchers (Chen et al., 2023; Sarthi et al., 2024) provide the text, candidate labels, and retrieve side information (e.g., descriptions of text and labels, or documents) for LLMs, and these LLMs output the labels of the text.
-Current work (Gao et al., 2023) demonstrates that this side information can provide context for the text and labels, enabling LLMs to comprehensively understand these inputs, which is a key component of the success of LLMs in text classification.
+An overview of GraphRAG (Edge et al., 2024).
+It first extracts and indexes entities and relations as graph nodes and edges.
+Then it performs Top-K similarity-based retrieval for query-related subgraphs.
+Finally, it classifies texts based on the subgraph summary and the query text.
+Recently, large language models (LLMs) (Sun et al., 2021; Jiang et al., 2023; Touvron et al., 2023; Achiam et al., 2023; Yang et al., 2024), pre-trained on extensive corpora, have achieved significant success due to their superior and comprehensive text understanding abilities.
+However, researchers (Gao et al., 2023; Peng et al., 2024) discover that LLMs can struggle to understand their inputs due to a lack of task-specific knowledge.
+Therefore, the Long Context RAG models (Chen et al., 2023; Sarthi et al., 2024) were proposed to combine the predicted text, candidate labels, and the retrieved side information (e.g., descriptions of text and labels, or documents) as LLM inputs, enable LLMs to understand the predicted text and labels more comprehensively with the help of side information from external sources.
 However, the incorporation of side information can further increase the input size and noise (Chen et al., 2023), which impedes the efficiency and effectiveness of LLMs (Sarthi et al., 2024).
-As a result, compression-based retrieval-augmented generation (RAG) approaches (Edge et al., 2024; Gutiérrez et al., 2024; Guo et al., 2024) are proposed to construct graphs by extracting key information from the side information sources.
-These models then only retrieve the information (e.g., graph path or subgraphs) relevant to the target text as the side information for LLMs, reducing the length of inputs for LLMs.
-However, existing compression-based RAG approaches (Edge et al., 2024; Guo et al., 2024; Gutiérrez et al., 2024) still have three limitations in the dynamic few-shot text classification task.
-Firstly, these approaches only extract information and build an information graph based on the side information of each text individually.
-In this way, they do not merge the information graphs of different text inputs into a single graph to obtain more comprehensive information for incoming target texts.
-Consequently, they fail to consider the correlation between side information of different text inputs, providing insufficient and incomplete context for LLMs.
-Secondly, existing approaches construct graphs by indexing extracted information linkages uniformly.
-They do not consider the varying importance and extraction confidence of each link, which may provide incorrect and unreliable context for LLMs.
-Thirdly, existing approaches select relevant information for each input text based on a globally predefined threshold.
-However, the optimal retrieval threshold can vary across different text samples, making the globally predefined threshold suboptimal for the entire dataset.
-To address the aforementioned issues, we propose a novel Graph-based Online Retrieval Augmented Generation framework for dynamic few-shot text classification, called GORAG.
-In general, GORAG constructs and maintains an adaptive online information graph by extracting side information from all target texts, tailoring graph retrieval for each input.
-Firstly, GORAG extracts keywords from the text using the LLM and links these keywords with the text’s ground truth label to represent the relationship between keywords and labels.
-Then, GORAG employs an edge-weighting mechanism when indexing edges into the information graph, assigning edge weights based on the importance of the keywords and their relevance to the respective text’s label.
-Secondly, when retrieving information from the graph for each text, GORAG retrieves candidate labels by constructing a minimum-cost spanning tree of text keywords mapped to the graph.
-After retrieval, the candidate labels serve as a filtered subset of the original target labels, which are then used to create input for the LLM.
-Since the generated spanning tree is determined solely by the graph information and the extracted text keywords, GORAG achieves adaptive retrieval without relying on any human-defined retrieval thresholds.
-We summarize the novel contributions as follows.
-  • We present a RAG framework for few-shot text classification tasks, namely GORAG.
-The GORAG framework consists of four stages: (1) Graph construction, where a graph is built with keywords from the training texts considering their correlation with the texts’ label; (2) Candidate label retrieval, where relevant labels are adaptively retrieved from the graph for label pre-filtering; (3) Prompt construction, where prompts are generated based on the filtered candidate labels and the testing texts; and (4) Online indexing, where the keywords of the un-labeled text are indexed to the constructed graph based on the prediction result.
-  • We develop a novel graph edge weighting mechanism based on the keyword nodes’ importance within the text corpus, which can be applied during both graph construction and online indexing.
-This mechanism enables our approach to effectively model the relevance between keywords and labels, thereby reducing the noise introduced to the graph.
-  • To avoid the need for human-selected thresholds during retrieval, we formulate the candidate label retrieval problem which akin to the NP-hard Steiner Tree problem.
-To solve this problem, we modified the Mehlhorn algorithm (Mehlhorn, 1988) which provides an efficient and effective solution to the problem to generate candidate types from the information graph.
-In the rest of this paper, we first present prelimnary and related work in section 2, the pipeline of our model GORAG in section 3, our experiments in section 4 and the final conclusion in section 5.
+As a result, Compression-based RAG approaches (Edge et al., 2024; Gutiérrez et al., 2024; Guo et al., 2024) are proposed to compress the length of context and extract the key information.
+Specifically, LongLLMLingua (Xia et al., 2021) and CompaAct (Yoon et al., 2024) propose filtering uninformative texts based on the query.
+However, these approaches may result in information loss and degrade overall performance.
+Then, to address this, graph RAG approaches (Edge et al., 2024; Guo et al., 2024; Gutiérrez et al., 2024; Peng et al., 2024; Zhang et al., 2025; Han et al., 2024a), such as GraphRAG (Edge et al., 2024) and LightRAG (Guo et al., 2024), propose to to index the unstructured texts into the structured graphs.
+For example, in Figure 2, the basic idea of graph RAG considers text entities as graph nodes, and text sentences describing relations between these entity nodes as graph edges, later, it retrieves subgraphs base on the query text, and generates short summaries of these subgraphs as the side information for LLMs, reducing the input length for LLMs.
+However, existing Compression RAG approaches (Edge et al., 2024; Guo et al., 2024; Gutiérrez et al., 2024) still have three issues in the DFSTC task.
+  1. (1)
+    Uniform-indexing issue: Some of these approaches construct graphs by indexing extracted text chunks uniformly.
+They do not consider the varying importance and extraction confidence of each text chunk, which may provide incorrect and unreliable context for LLMs.
+  2. (2)
+    Threshold-dependent issue: These approaches select relevant information for each input text based on a globally predefined threshold.
+However, the optimal retrieval threshold can vary across different text samples; hence, relying on a globally predefined threshold can be suboptimal for the entire dataset.
+  3. (3)
+    Narrow-source issue: These approaches only retrieve side information from the each few-shot training text and ignore the important information among different the query texts, consequently limiting their retrieve accuracy and comprehensiveness.
+To address these issues, we propose a novel hyperparameter-free Graph-based Online Retrieval Augmented Generation framework for dynamic few-shot text classification, called GORAG.
+In general, GORAG constructs and maintains an adaptive online weighted graph by extracting side information from all target texts, and tailoring the graph retrieval for each input.
+Firstly, GORAG constructs a weighted graph with keywords extracted from labeled texts as keyword nodes, and text labels as label nodes.
+The edge of this graph represents the relationship between keywords and labels.
+Secondly, GORAG employs an edge-weighting mechanism to assign different weights to edges created by the last step.
+An example of the constructed weighted graph is shown in Figure 3.
+The edge weights represent the keywords’ importance and relevance to the respective text’s label.
+Thirdly, GORAG constructs a minimum-cost spanning tree on the constructed weighted graph based on keywords of the query text, and then retrieves the label nodes within the spanning tree as candidate labels.
+Since the generated spanning tree is determined solely by the constructed graph and the keywords of the query text, GORAG achieves adaptive retrieval without relying on any human-defined retrieval thresholds, making the retrieval more precise regarding each query text.
+Lastly, GORAG applies an online indexing mechanism, which can enrich the weighted graph with the keywords from query texts, enhancing its few-shot learning performance by providing a more comprehensive retrieve source.
+We summarize our novel contributions as follows.
+  • We present a hyperparameter-free online RAG framework for DFSTC tasks, namely GORAG.
+The GORAG framework consists of three steps: graph construction, candidate label retrieval, and text classification, aim at addressing the uniform-index, threshold-dependent and narrow-source issues.
+  • We develop a novel graph edge weighting mechanism based on the text keywords’ importance within the text corpus, which enables our approach to effectively model the relevance between keywords and labels, thereby helping the more accurate retrieval.
+  • To avoid any human-selected thresholds, we formulate the candidate label retrieval problem which is akin to the NP-hard Steiner Tree problem.
+To solve this problem, provide an efficient and effective solution to generate candidate labels.
 2. Preliminary and Related Works
-In this section, we first introduce the preliminaries of dynamic few-shot text classification in Section 2.1 and then discuss the existing approaches in Section 2.2.
-The important notations used in this paper are listed in Table 1.
-2.1. Dynamic Few-shot Text Classification
+We first introduce the preliminaries of dynamic few-shot text classification in Section 2.1 and then discuss the related works for DFSTC task in Section 2.2.
+The important notations used in this paper are listed in Table 6 in Appendix A.
+2.1. Dynamic Few-shot Text Classification Task
 Text classification (Kowsari et al., 2019; Minaee et al., 2021; Li et al., 2022) is a key task in real-world application that involves assigning predefined labels $y\in\mathcal{Y}$ to text $t=(w_{1},\cdots,w_{|t|})$ based on its words $w_{i}$.
-It is widely applied in areas like sentiment analysis (Kouadri et al., 2020; maamar2022sa),query optimization (Yu et al., 2023; Yu and Litchfield, 2020) data integration (Dong and Rekatsinas, 2018), and schema matching (Sahay et al., 2020; Zhang et al., 2024a).
-Traditional approaches rely on large labeled datasets (Devlin, 2018; Liu, 2019), which may not always be available.
+Traditional approaches require fine-tuning with large labeled datasets (Devlin, 2018; Liu, 2019), which may not always be available.
 Recently, few-shot learning addresses this limitation by enabling models to classify text with only a small number of labeled examples per class (Meng et al., 2018; Meng et al., 2020).
-Dynamic classification (Parisi et al., 2019; Hanmo et al., 2024; Wang et al., 2024) introduces an additional challenge where new classes are introduced over multiple rounds, requiring the model to adapt to new classes while retaining knowledge of previously seen ones.
-Combining these aspects, Dynamic Few-Shot Text Classification (DFSTC) (Xia et al., 2021) allows the model to handle evolving classification tasks with minimal labeled data.
-In DFSTC, the model is provided with multiple rounds of new class updates.
-Specifically, In each round $r$, a new set of classes $\mathcal{Y}_{new}^{r}$ is introduced and the labelled dataset for $\mathcal{Y}_{\text{new}}^{r}$ is denoted as $\mathcal{D}_{new}^{r}=\cup_{y_{i}\in\mathcal{Y}_{\text{new}}^{r}}\{t_{j},y_{i}\}_{j=1}^{k}$, where per class $y\in\mathcal{Y}_{\text{new}}^{r}$ only has $k$ labeled examples $\{t_{j},y_{i}\}_{j=1}^{k}$.
-Also, we denote candidate cumulative labels from the first round to the $r$-th round as $\mathcal{Y}^{r}=\bigcup_{i=1}^{r}\mathcal{Y}_{\text{new}}^{i}$.
-Formally, at the $r$-th round, given the candidate labels $\mathcal{Y}^{r}$ and all labeled data $\mathcal{D}^{r}=\cup_{i=1}^{r}\mathcal{D}_{new}^{i}$, the target of DFSTC task is to learn a function $f^{r}_{\theta}$, which can learn scores for all target labels $\mathbf{s}^{r}_{k}={f}^{r}_{\theta}(t,\mathcal{Y}^{r})\in[0,1]^{|\mathcal{Y}^{r}|}$ for the unseen text $t_{k}$.
-Then, we can get the predicted label $y^{r*}_{k}\in\mathcal{Y}^{r}$ for the unseen text $t_{k}$ as follows.
-(1) $y^{r*}_{k}=\arg\max_{y^{r}_{k}\in\mathcal{Y}^{r}}({f}_{\theta}(t,\mathcal{Y}^{r})[y^{r}_{k}])$
-DFSTC is valuable for its ability to learn from limited labeled data, adapt to evolving class distributions, and address real-world scenarios where categories and data evolve over time.
-For example, in Round 1, the model might be trained on classes like Sports, Politics, and Technology.
-In Round 2, new classes such as Health and Education are introduced.
-The model must now classify texts into all five classes while retaining its knowledge of earlier ones.
-For instance, given the text ”The new vaccine shows promising results,” the model should classify it as Health.
-2.2. Dynamic Few-shot Text Classification Models
-Current dynamic few-shot text classification models can be broadly categorized into two types: Meta learning-based models and Data Augmentation-based models, and Large Language Model (LLM)-based models.
-2.2.1. Meta learning-based models
-Meta learning-based models train a meta-learner on base classes $C_{b}$ and then adapt it to new classes in each round.
-For instance, LEOPARD (Bansal et al., 2020) use the meta-learner to generate model parameters to predict classes in each round.
-ConEntail (Zhang et al., 2023a) trains the meta-learner to classify texts based on their nested entailments with class label names at each round.
-However, these models require a sufficient amount of training data to train the meta-learner, which differs from our experimental setting and limits their generalization ability.
-2.2.2. Data Augmentation-based models
-Data Augmentation-based models generate additional data contrastively based on the few-shot labeled data to train the classifier.
-For example, Entailment (Xia et al., 2021) creates extra data samples, by concatenating texts with class label names.
-WeSTClass (Meng et al., 2018) trains a document generator to produce pseudo documents contrastively based on few-shot training data.
-LOTClass (Meng et al., 2020) uses a Pretrained Language Model (PLM) to generate words semantically correlated with the class label name to help with the classification.
-However, due to the limited labeled data, the generated text data of these models can have very limited patterns, which makes them prone to overfitting (Liu et al., 2023).
+On the other hand, dynamic classification (Parisi et al., 2019; Hanmo et al., 2024; Wang et al., 2024d) introduces an additional challenge where new classes are introduced over multiple rounds, requiring the model to adapt to new classes while retaining knowledge of previously seen ones.
+Combining these aspects, Dynamic Few-Shot Text Classification (DFSTC) task allows the model to handle evolving classification tasks with minimal labeled data, and better suited for real world scnerio (Xia et al., 2021).
+In DFSTC task, the model is provided with multiple rounds of new class updates.
+Specifically, In each round $r$, a new set of classes $\mathcal{Y}_{new}^{r}$ is introduced, and the labelled dataset for $\mathcal{Y}_{new}^{r}$ is denoted as $\mathcal{D}_{new}^{r}=\cup_{y_{i}\in\mathcal{Y}_{new}^{r}}\{t_{j},y_{i}\}_{j=1}^{k}$, where per class $y_{i}\in\mathcal{Y}_{new}^{r}$ only has $k$ labeled examples $\{t_{j},y_{i}\}_{j=1}^{k}$.
+Also, we denote candidate cumulative labels from the first round to the $r$-th round as $\mathcal{Y}^{r}=\bigcup_{i=1}^{r}\mathcal{Y}_{new}^{i}$.
+Formally, at the $r$-th round, given the candidate labels $\mathcal{Y}^{r}$ and all labeled data $\mathcal{D}^{r}=\cup_{i=1}^{r}\mathcal{D}_{new}^{i}$, the target of DFSTC task is to learn a function $f^{r}_{\theta}$, which can learn scores for all target labels $\mathbf{s}^{r}={f}^{r}_{\theta}(t_{u},\mathcal{Y}^{r})\in[0,1]^{|\mathcal{Y}^{r}|}$ for the unseen text $t_{u}$.
+Then, we can get the predicted label $y^{r*}\in\mathcal{Y}^{r}$ for the unseen text $t_{u}$ as follows.
+(1) $y^{r*}=\arg\max_{y^{r}\in\mathcal{Y}^{r}}({f}_{\theta}^{r}(t_{u},\mathcal{Y}^{r})[y^{r}])$
+DFSTC is valuable for its ability to learn from limited labeled data, adapt to evolving class distributions, and address real-world scenarios where classes and data evolve over time.
+2.2. Dynamic Few-shot Text Classification Model
+Current DFSTC models can be broadly categorized into NN-based models and RAG-based models, RAG-based models can be further categorized into Long Context RAG models, and Compression-based RAG models.
+2.2.1. NN-based models
+NN-based models (Meng et al., 2018; Meng et al., 2020; Xia et al., 2021) generate additional data contrastively based on the few-shot labeled data and the text formed label names to train the NN-based classifier.
+However, due to the limited labeled data, the generated text data of these models can have very limited patterns, which makes them prone to overfitting (Liu et al., 2023; Liu et al., 2024c; Lei et al., 2023).
+Also, as the text-formed label names are not always available in real-world scenarios, these NN-based models may not always be applicable.
+Figure 3.
+An example of the constructed weighted graph, where red nodes denote the label nodes and black nodes denote the keyword nodes.
+The lower the edge cost is, the more related the keyword is to the respective label.
+Figure 4.
+An overview of GORAG in each round.
+In Part 1, GORAG constructs a weighted graph based on keywords extracted from the few-shot training data.
+In Part 2, GORAG performs adaptive graph retrieval and outputs the candidate labels, which is a subset of the original target label set.
+In Part 3, GORAG first classifies texts into one of the candidate labels, and then applies online indexing to update the graph with newly extracted nodes from query text.
+2.2.2. Long Context RAG models
+Recently, Large Language Model (LLM)-based models have undergone rapid development (Fernandez et al., 2023; Amer-Yahia et al., 2023; Zhang et al., 2024c; Miao et al., 2024; Snyder et al., 2024; Zhong et al., 2024; Yan et al., 2024; Chen et al., 2024b; Ding et al., 2024) and have been successfully adapted to various data mining tasks, including semantic parsing (Drozdov et al., 2022; Mekala et al., [n.d.]; Zhong et al., 2024; Lei et al., 2024; Sinha et al., 2024; Zhou et al., 2024), spatial-temporal mining (Zhang et al., 2024d; Li et al., 2024b; Yuan et al., 2024; Liu et al., 2024b; Wang et al., 2024a; Ning and Liu, 2024), and graph mining (Chen et al., 2024b; Tang et al., 2024; Cai et al., 2024; Jiang et al., 2024b; Cheng et al., 2024; Chen et al., 2024a; Fang et al., 2024).
+Notably, LLMs are inherently capable of inference without fine-tuning (Sun et al., 2021; Jiang et al., 2023; Touvron et al., 2023; Achiam et al., 2023; Yang et al., 2024), making them originally suitable for dynamic text classification tasks.
+However, the lack of fine-tuning can lead LLMs to generate incorrect answers, as they lack task-specific knowledge, these incorrect answers are often referred to as hallucinations (Zhang et al., 2023).
+To mitigate hallucinations, researchers try to retrieve text documents outside of LLM as side information to help the LLM inference, namely Long Context RAG models (Chen et al., 2023; Sarthi et al., 2024; Hu et al., 2024; Wu et al., 2024a; Cai et al., 2024; Oosterhuis et al., 2024; Yadav et al., 2024; Che et al., 2024).
+However, the retrieved contents from these models remain unstructured and can be lengthy, which impedes the efficiency and effectiveness of LLMs, leading to lost-in-the-middle issues (Liu et al., 2024a).
+2.2.3. Compression-based RAG models
+To address the issue suffered by Long Context RAG models, compression-based RAG models were proposed (Peng et al., 2024; Zhang et al., 2024b; Wu et al., 2024b; Li et al., 2024a; Zhang et al., 2024a; Jiang et al., 2024a; Guo et al., 2024; Gutiérrez et al., 2024), these models try to compress the side information to reduce the LLM input length.
+Based on how they compress model inputs, these models can be classified into prompt compressor models, and graph-based RAG models.
+On the one hand, prompt compressor models (Pan et al., 2024; Jiang et al., 2024a), apply LLM’s generation perplexity on to filter out un-important tokens in the model input.
+However, these models retrieve side information from each document individually, ignoring the inter-document correlations (Edge et al., 2024; Guo et al., 2024).
+On the other hand, graph-based RAG models (Edge et al., 2024; Guo et al., 2024; Gutiérrez et al., 2024) convert text chunks from the side information document as graph nodes, and the relation between these text chunks as graph edges.
+Based on the pre-defined retrieval threshold, they then retrieve a limited number of graph nodes and edges to represent the long document.
+However, the graph-based RAG models usually consider each graph edge uniformly and ignore the varying confidence and importance of relations between text chunks.
+Also, as the optimal retrieval threshold can vary across different data samples, their globally fixed threshold selected by humans may be suboptimal for the entire dataset.
+In this paper, we proposed a novel compression-based graph-RAG model, namely GORAG, which first constructs a weighted graph to model the inter-document correlations and varying confidence and importance of relations, then it achieves adaptive retrieval on this graph, avoiding any human-defined thresholds.
+3. Methodology
+3.1. Framework Overview
+To address the aforementioned uniform-indexing, threshold-dependent, and narrow-source issues, we propose GORAG, a novel approach that achieves adaptive retrieval by extracting valuable side information from a minimum-cost spanning tree generated on the constructed graph.
+As shown in Figure 4, GORAG consists of three core parts, i.e., index graph construction, graph retrieval, and text classification and online Index.
+Part 1: Graph Construction.
+It targets to construct the weighted graph $\mathcal{G}^{r}_{new}(\mathcal{V}^{r}_{new},\mathcal{E}^{r}_{new},\mathcal{W}^{r}_{new})$ based on the labeled data $\mathcal{D}_{new}^{r}=\cup_{y_{i}\in\mathcal{Y}_{\text{new}}^{r}}\{t_{j},y_{i}\}_{j=1}^{k}$ at the $r$-th round.
+An example the weighted graph created on WOS dataset is shown in Figure 3, graph formed like this will be used to provide retrieved-augmented information as context for query texts, enabling LLMs to better understand query texts and accurately predict their labels.
+Specifically, under each new data $\mathcal{D}_{new}^{r}$ at the $r$-th round, we first extract keywords $\mathcal{K}_{new}^{r}\in\{t_{j}\}_{j=1}^{|\mathcal{D}_{new}^{r}|}$ from the labeled training texts $\mathcal{T}^{r}_{new}=\{t_{j}\}_{j=1}^{|\mathcal{D}_{new}^{r}|}$.
+Then, we will assign an edge $e^{r}_{v,y}\in\mathcal{E}^{r}_{new}$ between each keyword $v\in\mathcal{K}_{new}^{r}$ and their respective text’s label $y\in\mathcal{Y}^{r}_{new}$.
+We then compute the weight $w^{r}_{v,y}\in\mathcal{W}^{r}_{new}$ for each edge $e^{r}_{v,y}\in\mathcal{E}^{r}_{new}$ based on the keyword’s importance with in the text corpus and relatedness to the label $y$.
+At last, we merge all graphs $\mathcal{G}^{r}_{new}(\mathcal{V}^{r}_{new},\mathcal{E}^{r}_{new},\mathcal{W}^{r}_{new})$ at each each round $r$ as the full graph $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r},\mathcal{W}^{r})$.
+More details can refer to Section 3.2.
+Part 2: Graph Retrieval.
+The Graph Retrieval process maps the extracted keyword nodes $\mathcal{V}^{t}$ from the query text $t$ to the constructed weighted graph $\mathcal{G}^{r}$ and generates a minimum-cost spanning tree on $\mathcal{G}^{r}$ that includes all these keywords.
+From this minimum-cost spanning tree (MST), the candidate label set $\hat{\mathcal{Y}}_{t}^{r}$ is obtained, which is a reduced subset of the original target label set $\mathcal{Y}^{r}$.
+As this MST solely depends on the graph itself and the extracted keywords, we can get rid of any human-defined thresholds and achieve adaptive retrieval.
+For more details, please refer to Section 3.3.
+Part 3: Text Classification and Online Index.
+In Text Classification part, we first classify the query text with a LLM.
+This process create LLM input with the query text $t$, the candidate labels $\hat{\mathcal{Y}}_{t}^{r}$ retrieved from the weighted graph $\mathcal{G}^{r}$, and descriptions $\mathcal{K}_{y_{i}}$ associated with each candidate label $y_{i}\in\hat{\mathcal{Y}}_{t}^{r}$ if text formed label names are available.
+The LLM will carry out classification with this input.
+After classification, the online indexing procedure dynamically indexes keywords, denoted as $\mathcal{V}_{\text{notexist}}^{t}$, which are extracted from the query text $t$ but are not in the existing graph $\mathcal{G}^{r}$.
+We then integrate these keywords into the graph to further enrich its structure, with edge weights assigned based on their importance within the text corpus and their relatedness to the predicted label $y^{*}_{t}$ for the query text $t$.
+It enhances the model’s ability to make more accurate predictions.
+For more details, please refer to Section 3.4.
+3.2. Part 1: Graph Construction
+In this subsection, we introduce the graph construction procedure of GORAG, which includes a novel edge-weighting mechanism to address the uniform indexing issue.
+The pseudo code of graph construction is shown in Algorithm 1 in Appendix Appendix C.
+For each round $r$, given its respective labeled training texts $\mathcal{D}_{new}^{r}=\cup_{y_{i}\in\mathcal{Y}_{\text{new}}^{r}}\{t_{j},y_{i}\}_{j=1}^{k}$, GORAG first extracts text keywords $\mathcal{K}^{r}_{new}$ from $\mathcal{D}_{new}^{r}$.
+Specifically, it uses the LLM model with an extraction instruction prompt $p_{extract}=$Please extract some keywords from the following passage” for this extaction.
+Then, we can get all text keywords $\mathcal{K}^{r}_{new}$ based on texts $\mathcal{D}^{r}_{new}$ as follows.
+(2) $\mathcal{K}^{r}_{new}=\bigcup_{t\in\mathcal{D}^{r}_{new}}LLM(p_{extract},t).\vspace{-5px}$
+These keywords are served as graph nodes, we denote them as keyword nodes.
+Also, GORAG incorporats new candidate labels $\mathcal{Y}^{r}_{new}$ at the $r$-th round as graph nodes, denoted as label nodes.
+Hence, the graph node set $\mathcal{V}^{r}_{new}$ at the $r$-th round can be obtained as follows.
+(3) $\mathcal{V}_{new}^{r}=\mathcal{K}_{new}^{r}\cup\mathcal{Y}_{new}^{r}.$
+Then, GORAG add edges between each keyword node $v\in\mathcal{K}_{{new}}^{r}$ to its corresponding label node $y\in\mathcal{Y}_{{new}}^{r}$, indicating that the keyword $v$ appears in texts associated with the label $y$.
+Next, considering the keywords $\mathcal{K}_{new}^{r}$ from the labeled text are not uniformly related to the text’s label $\mathcal{Y}_{new}^{r}$, we apply an edge weighting mechanism to assign a weight $w_{v,y}$ to each keyword-label edge.
+This weight can be regarded as the correlation between keywords and each label.
+Firstly, we apply the normalized TF-IDF score (Sammut and Webb, 2010) as our correlation score $CS(v,t)$ to measure the importance and relatedness of a particular keyword $v\in\mathcal{K}_{new}^{r}$ w.r.t.
+the respective text $t$ it extracted from:
+(4) $\vspace{-5px}CS(v,t)=\frac{count(v,t)}{|t|}\times log\frac{|\mathcal{T}^{r}|}{1+|t_{j}:v\in t_{j},\ t_{j}\in\mathcal{T}^{r}|},$
+where $\mathcal{T}^{r}=\cup_{i=1}^{r}\mathcal{T}_{new}^{i}$ denotes all training and query texts seen so far, $count(v,t)$ is the number of times that the term $v$ appears in the text $t$, and $|t_{j}:v\in t_{j},\ t_{j}\in\mathcal{T}^{r}|$ denotes the number of texts in the corpus $\mathcal{T}^{r}$ that contain the keyword $v$.
+As the keyword $v$ can be extracted from multiple texts with different labels in different rounds, the final edge weight $w_{v,y}^{r}$ of edge $e_{v,y}$ at the $r$-th round is calculated as an average of all weights from all seen texts with label $y$:
+(5) $\vspace{-3px}w_{v,y}^{r}=\frac{\sum_{t_{j}\in\mathcal{T}^{r,y,v}}1-CS(v,t_{j})}{|\mathcal{T}^{r,y,v}|},$
+where $\mathcal{T}^{r,y,v}=\{t_{j}|v\in t_{j}\land t_{j}\in\mathcal{T}^{r,y}\}$ is the text that contains keyword $v$ and labeled $y$.
+We denote the generated weighted graph for $\mathcal{K}_{new}^{r}$ and $\mathcal{Y}_{new}^{r}$ at the $r$-th round as $\mathcal{G}^{r}_{new}(\mathcal{V}_{new}^{r},\mathcal{E}_{new}^{r},\mathcal{W}_{new}^{r})$, where $\mathcal{V}^{r}_{new}$, $\mathcal{E}^{r}_{new}$, and $\mathcal{W}^{r}_{new}$ denotes the node, edge and edge weight set respectively, an example of the generated graph is shown in Figure 3.
+Lastly, $\mathcal{G}^{r}_{new}$ will be merged into the graph from previous rounds $\mathcal{G}^{r-1}$ to form the full graph $\mathcal{G}^{r}$ at round $r$:
+(6) $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r},\mathcal{W}^{r}),$
+$\mathcal{V}^{r}=\mathcal{V}^{r}_{new}\cup\mathcal{V}^{r-1},$ $\ \mathcal{E}^{r}=\mathcal{E}^{r}_{new}\cup\mathcal{E}^{r-1},\ \mathcal{W}^{r}=\mathcal{W}^{r}_{new}\cup\mathcal{W}^{r-1}.$
+Particularly, we define $\mathcal{G}^{0}=\emptyset$ and $\ r\geq 1$.
+Also, to guarantee the graph connectivity of the resulting graph, we add edges between every new label node $y_{n}\in\mathcal{Y}_{new}^{r}$ and each old label node $y_{o}\in\mathcal{Y}^{r-1}$, the edge weight $w_{y_{n},y_{o}}^{r}$ of edge $e_{y_{n},y_{o}}$ at round $r$ is defined as the weighted average of all edge weights that link keyword nodes with label node $y_{n}$ or $y_{o}$:
+(7) $\vspace{-5px}\mathcal{M}_{y}^{r}=\{v\ |\ v\in\mathcal{N}^{r}(y)\ \land\ v\notin\mathcal{Y}^{r}\}\vspace{-5px}$
+(8) $\vspace{-3px}w_{y_{n},y_{o}}^{r}=\frac{1}{2}\big{(}\frac{\sum_{v\in\mathcal{M}_{y_{n}}^{r}}w_{v,y_{n}}^{r}}{2\times|\mathcal{M}_{y_{n}}^{r}|}+\frac{\sum_{v\in\mathcal{M}_{y_{o}}^{r}}w_{v,y_{o}}^{r}}{2\times|\mathcal{M}_{y_{o}}^{r}|}\big{)}$
+where $\mathcal{N}^{r}(y)$ denote the neighbor nodes of label node $y$ in $\mathcal{G}^{r}$.
+After merge, the graph $\mathcal{G}^{r}$ would be used for future retrieval, and be further updated by GORAG’s online indexing mechanism.
+3.3. Part 2: Graph Retrieval
+In this subsection, we introduce the graph retrieval procedure of GORAG.
+With the graph constructed in Part 1, GORAG can adaptively retrieve a set of candidate class labels with keywords extracted from query texts without any human-defined thresholds, addressing the threshold dependent issue.
+To begin with, GORAG extract keywords $\mathcal{V}_{test}^{t}$ for each query text $t\in\mathcal{T}_{test}$ in the same manner with Equation (2), then, $\mathcal{V}_{test}^{t}$ would be splitted into two subsets:
+(9) $\mathcal{V}_{exist}^{t}\cup\mathcal{V}_{notexist}^{t}=\mathcal{V}_{test}^{t},\ \mathcal{V}_{exist}^{t}\cap\mathcal{V}_{notexist}^{t}=\emptyset,$
+where $\mathcal{V}_{exist}^{t}$ and $\mathcal{V}_{notexist}^{t}$ denotes the keywords in $\mathcal{V}_{test}^{t}$ that already exist and not yet exist in $\mathcal{G}^{r}$ at current round $r$ respectively.
+Later, $\mathcal{V}_{exist}^{t}$ would be applied for achieving adaptive retrieval, and $\mathcal{V}_{notexist}^{t}$ would be applied for online indexing to further enrich the constructed graph (Further illustrated in Section 3.4).
+To achieve the adaptive retrieval, GORAG trys to find the minimum cost spanning tree (MST) that contain all keyword nodes within $\mathcal{V}_{exist}^{t}$, and then retrieve all label nodes within the generated MST as candidate labels.
+The intuition behind this approach is that a MST spans the entire graph to cover all given nodes with the smallest possible spanning cost.
+Consequently, label nodes within the generated MST can be considered important for demonstrating the features of the given keyword node set.
+As the generation of an MST is a classical combinatorial optimization problem with an optimal solution determined solely by the set of given nodes.
+By generating the MST and retrieving all label nodes within it, we eliminate the need for any human-defined thresholds, achieving adaptive retrieval.
+Table 1.
+Statistics of the WOS, Reuters, and IFS-Rel datasets, where we divide their original labels into multiple rounds.
+We achieve a balanced testing set on the Reuters and IFS-Rel datasets by assigning each label 10 and 40 testing samples, respectively.
+Dataset Avg. Text Token R1 R2 R3 R4 Total
+Testing data Label # Testing data Label # Testing data Label # Testing data Label # Testing data
+WOS (Kowsari et al., 2017) 200 2,417 32 2,842 53 2,251 30 1,886 18 9,396
+Reuters (Bao et al., 2020) 168 80 8 80 8 80 8 70 7 310
+IFS-Rel (Xia et al., 2021) 105 640 16 640 16 640 16 640 16 2,560
+Definition 3.1 (Adaptive Candidate label Generation Problem).
+Given an undirected weighted connected graph $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r},\mathcal{W}^{r})$, a set of keywords $\mathcal{V}_{exist}^{t}$ extracted from text $t$ that can be mapped to nodes in $\mathcal{V}^{r}$, and the target label set $\mathcal{Y}^{r}\in\mathcal{V}^{r}$ at the $r$-th round, our target is to find a set of labels nodes ${\mathcal{Y}}^{r}_{t}\in{\mathcal{Y}}^{r}$.
+Firstly, we identify a subgraph $\mathcal{G}^{r}_{t}(\mathcal{V}^{r}_{t},\mathcal{E}^{r}_{t},\mathcal{W}^{r}_{t})$ of $\mathcal{G}^{r}$ by minimizing the edge weight sum as follows.
+$\min_{e^{r}_{u.v}\in\mathcal{E}^{r}_{t}}w^{r}_{u,v}$
+(10) $s.t.$ $v\in\mathcal{V}^{r}_{t},\forall v\in\mathcal{V}_{exist}^{t}$
+Then, since the subgraph nodes $\mathcal{V}^{r}_{t}$ contains both keyword nodes and label nodes, we take the label nodes ${\mathcal{Y}}^{r}_{t}\in\mathcal{V}^{r}_{t}$ in the gernerated sub-graph as our target candidate nodes for the text $t$.
+Since this problem is the NP-hard (Su et al., 2020), it is infeasible to obtain the optimal result in polynomial time.
+Therefore, to solve this problem, we propose a greedy algorithm based on the Melhorn’s algorithm (Mehlhorn, 1988a).
+Our algorithm generates the candidate label set ${\mathcal{Y}}^{r}_{t}\subseteq\mathcal{Y}^{r}$ for text $t$ at the $r$-th round.
+The pseudo code of our algorithm is shown in Appendix C, following the Melhorn’s algorithm, our algorithm can also achieve 2-approximate.
+Firstly, we calculate the shortest path between each keyword node $v$ in $\mathcal{V}_{exist}^{t}$ to all other nodes in $\mathcal{G}^{r}$ by calculating the minimum spanning tree $MST$ of $\mathcal{G}^{r}$ Here, the keyword nodes $v$ in $\mathcal{V}_{exist}^{t}$ are served as terminal nodes that determines the final generated candidate labels w.r.t.
+to the weighted graph $\mathcal{G}^{r}$ Secondly, we create a new auxiliary graph $H$ where the edges represent the shortest paths between the closest terminal nodes.
+Thirdly, we construct the minimum spanning tree $MST^{\prime}$ of the auxiliary graph $H$ , and then add the shortest paths between each two nodes in $MST^{\prime}$ to the Steiner Tree $ST$.
+Lastly our candidate labels $\hat{\mathcal{Y}}^{r}_{t}$ for text $t$ are calculated as the interselect of $ST$ and all target labels $\mathcal{Y}^{r}$ at round $r$: $ST\cap\mathcal{Y}^{r}$.
+Please refer to Appendix C for more details.
+3.4. Part 3: Text Classification and Online Index
+In this subsection, we introduce GORAG’s classification part, where it futher enrich the weighted graph by adding query text information with the online indexing mechanism, addressing the narrow source issue.
+To begin with, we introduce how GORAG performs text classification based the LLM.
+Specifically, given each unlabled query text $t$, GORAG utilizes LLM to predict its class label $y^{*}$ by constructing an input prompt $c_{t}$ as follows:
+(11) $c_{t}=Concat(t,\mathcal{K}^{t},\hat{\mathcal{Y}}^{r}_{t},\mathcal{K}_{\hat{\mathcal{Y}}^{r}_{t}}).$
+$c_{t}$ can be considered as the text concatenation of the extracted keywords $\mathcal{K}^{t}$ from the text $t$, and the candidate labels $\hat{\mathcal{Y}}^{r}_{t}$ obtained by Algorithm 2.
+Also, the $\mathcal{K}_{\hat{\mathcal{Y}}^{r}_{t}}=\{\mathcal{K}_{y_{i}}\}_{i=1}^{|\hat{\mathcal{Y}}^{r}_{t}|}$ and $\mathcal{K}_{y_{i}}$ are the representative keywords of each label $y_{i}\in\hat{\mathcal{Y}}^{r}_{t}$.
+(12) $\mathcal{K}_{y_{i}}=LLM(p_{gen},\mathcal{D}^{r}),$
+which were generated by LLM with the label’s semantic label name if avaiable. Next, with a classification instruction prompt $p_{classify}$.
+(13) $\vspace{-5px}y^{*}_{t}=LLM(p_{classify},c_{t}),$
+the LLM would try to select the best-suited label $y^{*}_{t}\in\mathcal{Y}^{t}$ to annotate the text $t$.
+To fully leverage the text-extracted keywords, GORAG utilizes an online indexing mechanism to incrementally update keywords that do not yet exist in the weighted graph $\mathcal{G}^{r}$ at the $r$-th round to $\mathcal{G}^{r}$ based on the text $t$’s predicted label $y^{*}_{t}$.
+To be specific, each keyword node $v\in\mathcal{V}_{notexist}^{t}$ would be added to the original graph’s node set $\mathcal{V}^{r}$ and be assigned with an edge $e_{v,y}$ connecting it with the text $t$’s predicted label $y_{t}^{*}$:
+(14) $\vspace{-5px}\mathcal{V}^{r}=\mathcal{V}^{r}\cup\mathcal{V}_{notexist}^{t},\ \mathcal{E}^{r}=\mathcal{E}^{r}\cup\mathcal{E}_{oi}^{t},$
+where $\mathcal{E}_{oi}^{t}=\{e_{v,y_{t}^{*}}|v\in\mathcal{V}_{notexist}^{t}\}$ denotes the set of all newly assigned edges $e_{v,y}$ between keyword node $v$ and its predicted label $y_{t}^{*}$, for these newly assigned edges, their weight is calculated with the edge weighting mechanism illustrated in Equation (5).
+This online indexing mechanism has two purposes.
+Firstly, it incrementally enriches the weighted graph $\mathcal{G}^{r}$ by incorporating new, previously unseen keywords, thus expanding the graph’s vocabulary and improving its adaptability.
+Secondly, by linking these new keywords to their predicted labels with weighted edges, this mechanism captures their relevance and context more effectively, enabling better utilization of the graph for future predictions.
+4. Experiments
+In this section, we present the experimental evaluation of our framework GORAG on the DFSTC task.
+We compare the performance of GORAG against six effective baselines spanning three technical categories, using three datasets with distinct characteristics.
+Specifically, we first outline the experimental setup in Section 4.1, including details on datasets, and evaluation metrics.
+Due to the space limitation, we put the details of baselines, and hyperparameter configurations in Appendix Section E.
+Next, we report the experimental results, focusing on both effectiveness and efficiency evaluations, in Section 4.2.
+Finally, we conduct an ablation study and a case study, presented in Section 4.5 and Appendix Section E.5.
+4.1. Experiment Settings
+4.1.1. Few-shot setting
+In this paper, we employ 1-shot, 5-shot, and 10-shot settings for few-shot training, where each setting corresponds to using 1, 5, and 10 labeled training samples per class, respectively.
+We omit experiments with more than 10 labeled samples per class, as for WOS dataset, there are already over 1300 labeled training data under 10-shot setting.
+4.1.2. Datasets
+We select the Web-Of-Science (WOS) dataset (Kowsari et al., 2017) and the Reuters few-shot text classification dataset by (Bao et al., 2020) for evaluating GORAG’s performance.
+We chose these datasets because they have a considerable amount of classes and text tokens, which is complex and can reflect the real-world scenario.
+As shown in Table 1, texts from these datasets can contain multiple sentences and have average tokens ranging from 105 to 200 per text.
+The IFS-Rel dataset was extended from its original version in (Xia et al., 2021), which only has 32 labels in total and an average of 25 tokens per text.
+Table 2. Classification accuracy on WOS, Reuters, and IFS-Rel dataset.
+Dataset Category Model Round
+R1 R2 R3 R4
+1-shot 5-shot 10-shot 1-shot 5-shot 10-shot 1-shot 5-shot 10-shot 1-shot 5-shot 10-shot
+WOS NN-based Entailment (Xia et al., 2021) 0.3695 0.3823 0.4187 0.3994 0.4471 0.4222 0.4510 0.4857 0.4787 0.4030 0.4387 0.4442
+Long Context RAG NaiveRAG (Gao et al., 2023) 0.3885 0.3904 0.3897 0.2267 0.2154 0.2187 0.1821 0.1475 0.1799 0.1653 0.1556 0.1649
+Propositionizer (Chen et al., 2023) 0.1241 0.1306 0.1297 0.1074 0.1421 0.1303 0.1771 0.1645 0.1603 0.1611 0.1637 0.1656
+Compression RAG LongLLMLingua (Jiang et al., 2024a) 0.3806 0.3823 0.3901 0.2155 0.2202 0.2198 0.1770 0.1567 0.1608 0.1468 0.1382 0.1493
+GraphRAG (Edge et al., 2024) 0.3852 0.3897 0.3906 0.2213 0.2197 0.2219 0.1816 0.1770 0.1786 0.1641 0.1634 0.1625
+LightRAG (Guo et al., 2024) 0.3930 0.3806 0.3815 0.2202 0.2216 0.2145 0.1743 0.1767 0.1799 0.1625 0.1626 0.1632
+GORAG 0.4866 0.4990 0.5134 0.4790 0.5109 0.5284 0.4736 0.4996 0.5234 0.4380 0.4717 0.4929
+Reuters NN-based Entailment (Xia et al., 2021) 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
+Long Context RAG NaiveRAG (Gao et al., 2023) 0.0000 0.0000 0.0375 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
+Propositionizer (Chen et al., 2023) 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
+Compression-RAG LongLLMLingua (Jiang et al., 2024a) 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0258 0.0065 0.0085
+GraphRAG (Edge et al., 2024) 0.1375 0.1625 0.1000 0.0688 0.0813 0.0500 0.0375 0.0417 0.0417 0.0291 0.0375 0.0375
+LightRAG (Guo et al., 2024) 0.0500 0.1375 0.1125 0.0250 0.0813 0.0563 0.0333 0.0333 0.0333 0.0125 0.0333 0.0208
+GORAG 0.0875 0.0875 0.1000 0.1750 0.1688 0.1438 0.1667 0.1958 0.2167 0.1667 0.2516 0.2193
+IFS-Rel NN-based Entailment (Xia et al., 2021) 0.3391 0.6046 0.5859 0.4008 0.5516 0.5828 0.3061 0.4703 0.5193 0.2971 0.4395 0.4895
+Long Context RAG NaiveRAG (Gao et al., 2023) 0.4844 0.4719 0.4828 0.3711 0.3555 0.3656 0.2708 0.2750 0.2625 0.2516 0.2465 0.2527
+Propositionizer (Chen et al., 2023) 0.1719 0.1828 0.1938 0.1602 0.1727 0.1750 0.1792 0.1552 0.1770 0.1574 0.1688 0.1664
+Compression-RAG LongLLMLingua (Jiang et al., 2024a) 0.4703 0.5203 0.4297 0.3458 0.3055 0.3733 0.2452 0.2206 0.2713 0.2154 0.2029 0.1997
+GraphRAG (Edge et al., 2024) 0.4797 0.4656 0.4734 0.3672 0.3641 0.3641 0.2740 0.2776 0.2708 0.2453 0.2465 0.2426
+LightRAG (Guo et al., 2024) 0.4484 0.4234 0.4656 0.3531 0.3477 0.3523 0.2760 0.2760 0.2682 0.2340 0.2367 0.2391
+GORAG 0.4875 0.5266 0.5313 0.4195 0.3781 0.4374 0.2953 0.2974 0.3328 0.2973 0.2571 0.3034
+Table 3. The size of the constructed weighted graph in each round.
+Dataset Model Round
+R1 R2 R3 R4 After R4
+Node # Edge # Node # Edge # Node # Edge # Node # Edge # Node # Edge #
+WOS GORAG offline 1,149 1,405 3,681 4,587 4,947 6,204 5,438 6,879 5,438 6,879
+GORAG 1,149 1,405 12,108 12,245 25,357 26,290 34,521 35,919 44,283 45,973
+Reuters GORAG offline 117 119 192 200 263 284 331 366 331 366
+GORAG 117 119 584 594 1,013 1,046 1,342 1,405 1,533 1,600
+IFS-Rel GORAG offline 189 213 242 275 372 469 454 594 454 594
+GORAG 189 213 5,299 5,503 9,417 9,934 13,293 14,074 16,705 17,486
+We split the WOS dataset into 4, 6 and 8 rounds, the Reuters and IFS-Rel datasets into 4 rounds, in order to maintain a comparable number of new label and testing data in each round.
+Due to the space limitation, we mainly display our experinents on the 4 rounds spit version of both datasets in Table 2, and the experiment result of 6 and 8 rounds are shown in Figure 5(a) and 5(b).
+Different from the WOS and IFS-Rel datasets, the labels of Reuters are in numeric formats, and their text-formed label names are not available.
+According to Table 2, this feature makes many baselines fail to carry out classification task on the Reuters dataset.
+Further statistics of these two dataset being splitted into 4 rounds are shown in Table 1.
+4.1.3. Evaluation Metrics
+In this paper, aligning with previous DFSTC models (Xia et al., 2021; Meng et al., 2020), we use classification accuracy as the main evaluation metric.
+Given that LLMs can generate arbitrary outputs that may not precisely match the provided labels, we consider a classification correct only if the LLM’s output exactly matches the ground-truth label name or label number, if the model cannot give out a class within the provided target label set, we denote this classification as a hallucination.
+We further provide the macro recall score and the hallucination rate on the most complicated and imbalanced test dataset WOS in Table 8 in Appendix.
+4.1.4. Baselines
+As shown in Table 2, we compare GORAG’s performance with 7 baselines from 3 technical categories for few-shot experiments, they are NN-based Entailment (Xia et al., 2021), Long Context-based NaiveRAG (Gao et al., 2023), Propositionizer (Chen et al., 2023), and Compression-based LongLLMLingua (Jiang et al., 2024a), GraphRAG (Edge et al., 2024), and LightRAG (Guo et al., 2024).
+Further illustrations of these chosen baselines are listed in subsection E.1 in Appendix.
+For GraphRAG and NaiveRAG, we use the implementation from (nan, [n.d.]) which optimizes their original code and achieves better time efficiency while not affect the performance; For all other baselines, we use their official implementations.
+Also, long text inputs affect LLM performance and increase costs (Chen et al., 2023; Jiang et al., 2024a; Liu et al., 2024a).
+WOS and Reuters datasets, with lengthy texts and many labels, caused GPU out-of-memory (e.g., llama3-8B, 1-shot, A100-40GB).
+Thus, we compare open-source LLMs with GORAG in the 0-shot setting (Table 4).
+4.2. Few-shot Experiments
+4.2.1. Effectiveness Evaluation
+As shown in Table 2, GORAG achieves the best classification accuracy over all four rounds with 1-shot and 5-shot labeled indexing data on WOS dataset, it surpasses all RAG-based baselines as well as the state-of-the-art model Entailment.
+Compared with Entailment, when apply 1-shot setting, GORAG achieves at most 31.6% accuracy gain at the first round.
+On Reuters dataset, GORAG achieves the best classification accuracy over the last 3 rounds with 1-shot indexing data.
+Furthermore, based on the experiment results in Table 2, we can have following observations:
+  • Firstly, NN-based Entailment fails on Reuters due to its need for text-formed label names.
+As noted in prior studies (Liu et al., 2023; Liu et al., 2024c; Lei et al., 2023), it overfits easily, with recall and hallucination issues shown in Table 8.
+Further analysis is in Appendix subsection E.3.
+  • Secondly NaiveRAG performs better than Entailment in round 1 of the WOS and IFS-Rel datasets by avoiding overfitting, but its long LLM inputs hinder its ability to adapt to dynamic label updates, leading to a significant performance drop from rounds 2 to 4.
+While Propositionizer retrieves more fine-grained results, its lack of task awareness results in consistently poor performance on DFSTC tasks.
+On the Reuters dataset, both Long Context RAG baselines fail when label names are unavailable, as the absence of semantic-rich labels prevents effective classification, especially with long LLM inputs.
+  • Thirdly, compression-based RAG models shorten LLM inputs and emphasize labels and key terms but only slightly improve performance across datasets.
+They struggle with dynamic label updates due to issues like uniform indexing, threshold dependency, and narrow sources, leading to inaccurate retrievals.
+Additionally, failing to compress label sets results in lengthy LLM inputs as new labels are introduced in each round.
+  • Lastly, GORAG addresses the issues of Compression-based RAG models.
+From R1 to R4, GORAG adapts better to dynamic label updates.
+On the Reuters dataset, when label names are not available, GORAG achieves comparable or better performance than other baselines.
+According to Table 3, a vast amount of knowledge is added to the weighted graph by the online indexing mechanism, and contributes to our effectiveness.
+4.2.2. Efficiency Evaluation
+To evaluate the efficiency of GORAG, we compare GORAG’s indexing time cost with other baselines’ training or indexing time cost on the 4 round split version of WOS and Reuters datasets, the results are shown in Figure 5(c) and 5(d).
+For both datasets, Propositionizer achieves the worst indexing efficiency.
+This is because it employs an additional fine-tuned LLM to convert each side information into propositions.
+In this process, this extra LLM rewrite each labeled or query text and generates a long list of propositions, introducing more efficiency overhead compared to other models that use the LLM solely for classification or keyword extraction, which typically generate only a few tokens.
+It is also worth noting that, although GORAG’s online indexing mechanism introduces a vast amount of knowledge into the weighted graph, we can still achieve comparable or better time efficiency with other baselines.
+(a) 1-shot setting.
+(b) 5-shot setting.
+(c) Time cost on WOS dataset.
+(d) Time cost on Reuters dataset.
+Figure 5. Experiment result of WOS dataset under 1-shot and 5-shot settings for at most 8 rounds, and model Training/Indexing time cost on two datasets under 1-shot setting.
+4.3. Zero-shot Experiments
+To further study GORAG’s 0-shot ability, we compare GORAG’s performance with some widely applied open-source LLM models without RAG approaches.
+We conducted experiments on the WOS dataset by providing only the label names, without any labeled data to GORAG for creating the graph, the result is shown in Table 4.
+In the 0-shot setting, GORAG first generates label descriptions based on the label names, then extracts keywords from these descriptions, without using any information from labeled texts.
+Note that the Entailment model cannot be applied in this setting, as they require labeled data for training.
+For other RAG-based baselines, they are considered equivalent to their backbone LLMs in the 0-shot setting due to a lack of external information sources.
+As shown by the results, GORAG achieves comparable performance with other open-source LLM models in the first round and outperforms them in all subsequent rounds.
+This is because, when the number of target labels is small, the benefit of compressing the target label set is not as pronounced.
+However, as the number of labels increases, the importance of filtering the label set becomes more significant, allowing GORAG to consistently outperform other open-source LLM models.
+Table 4. Zero-shot experiment result on WOS dataset.
+Model Round
+R1 R2 R3 R4
+Qwen2-7B (Yang et al., 2024) 0.3322 0.1790 0.1451 0.1444
+Mistral0.3-7B (Jiang et al., 2023) 0.1436 0.0540 0.0336 0.0270
+LLaMA3-8B (Touvron et al., 2023) 0.3351 0.1614 0.1161 0.0930
+GORAG 0.3305 0.2567 0.2230 0.2102
+4.4. Generalization Evaluation
+To evaluate the generalization ability of GORAG, we conducted additional experiments on the 6 and 8 round split versions of the WOS dataset under 1-shot and 5-shot settings, as the WOS dataset contains more testing data and labels than Reuters dataset, and more complex texts than the IFS-Rel dataset.
+In Figure 5, the results for rounds 1 to 4 are obtained from the 4-round split version, rounds 5 and 6 are obtained from the 6-round split version, and rounds 7 and 8 are obtained from the 8-round split version of the WOS dataset.
+This setting can ensure a sufficient amount of testing data at each round, As illustrated in Figure 5, the classification accuracy of GraphRAG and NaiveRAG drops significantly from round 1 to 8, while GORAG maintains competitive classification accuracy as the number of rounds increases through 1 to 8, with both 1 and 5-shot setting, demonstrating the importance of a more precise retrieval.
+Table 5. Ablation experiments on WOS for GORAG’s variants.
+Dataset Model Round
+R1 R2 R3 R4
+WOS GORAG offline 0.3063 0.2302 0.2455 0.2156
+GORAG unit 0.4706 0.4394 0.4407 0.3899
+GORAG keyword 0.4746 0.4606 0.4455 0.4030
+GORAG 0.4862 0.4649 0.4814 0.4210
+Reuters GORAG offline 0.0000 0.1000 0.0917 0.1032
+GORAG unit 0.0000 0.0625 0.1667 0.0795
+GORAG keyword 0.0000 0.1000 0.1167 0.1438
+GORAG 0.0875 0.1750 0.1667 0.1667
+4.5. Ablation Study
+To analyze GORAG’s superior performance, we conduct an extensive ablation study in the 1-shot setting, as RAG models show minimal performance changes across shot settings.
+Specifically, we examine how different GORAG components impact classification performance using its variants.
+GORAG offline removes the online indexing mechanism.
+GORAG unit removes the edge weighting mechanism, every edge in this variant is assigned with weight 1.
+GORAG keyword only uses the keyword extracted from the text to create LLM classification input, rather than the whole text.
+As shown in Table 5, firstly, GORAG offline achieves the worst accuracy, this is because it lacks a comprehensive retrieval source, making its generated candidate labels inaccurate, hence limits its performance.
+Secondly, GORAG unit outperforms GORAG offline, but still sub-optimal to GORAG, demonstrates the importance of modeling the varying importance between keywords and each label, which leads to accurate retrievals.
+Lastly, GORAG keywords demonstrates the important information for text classification are mostly text keywords, as it achieves comparable or better performance than other ablated baselines.
+5. Conclusion
+In this paper, we propose GORAG, a Graph-based Online Retrieval Augmented Generation framework for the Dynamic Few-shot Text Classification (DFSTC) task.
+Extensive experiments on three datasets with different characteristics demonstrate the effectiveness of GORAG in classifying texts with only a limited number or even no labeled data.
+Additionally, GORAG shows its effectiveness in adapting to the dynamic updates of target labels by adaptively retrieving candidate labels to filter the large target label set in each update round.
+With extensive of ablation studies, we confirm that GORAG’s edge weighting, adaptive retrieval, and online indexing mechanisms contribute to its effectiveness.
+For future work, we aim to further enhance GORAG’s performance and explore its application in more general scenarios.
+Appendix A Notations
+Due to the space limitation, we put the summary table on important notations here.
 Table 6. Summary on Important Notations.
@@
 $f^{r}_{\theta}$ The function that learn all target labels’ score for texts.
-$\mathbf{s}^{r}_{k}$ The score for all target labels at round $r$.
-$y^{r*}_{k}$ The predicted label for the text.
-$\mathcal{G}^{r}(\mathcal{V},\mathcal{E})$ Information graph with node set $\mathcal{V}$ and edge set $\mathcal{E}.$
+$\mathbf{s}^{r}$ The score for all target labels at round $r$.
+$y^{r*}$ The predicted label for the text at round $r$.
+$\mathcal{G}^{r}(\mathcal{V},\mathcal{E})$ weighted graph with node set $\mathcal{V}$ and edge set $\mathcal{E}.$
 $\mathcal{Y}^{r}$, $\mathcal{Y}^{r}_{new}$ All target label and new target labels at round $r$.
-$\mathcal{D}^{r}$, $\mathcal{D}^{r}_{new}$ The labeled text for all labels and new labels at round $r$.
-$\mathcal{T}^{r}$, $\mathcal{T}_{new}^{r}$ The text of all labels and new labels at round $r$.
-$\mathcal{T}_{test}^{r}$ The testing text for at round $r$.
+$\mathcal{D}^{r}$, $\mathcal{D}^{r}_{new}$ The labeled training text for all seen labels at round $r$.
+$\mathcal{T}^{r}$, $\mathcal{T}_{new}^{r}$ All seen texts at round $r$.
+$\mathcal{T}_{test}^{r}$ The query text for at round $r$.
 $\mathcal{K}_{new}^{r}$ The extracted keyword set for new labels of round $r$.
 $\mathcal{V}_{new}^{r}$ The set of new graph nodes to be added in round $r$.
-$\mathcal{N}(\cdot)$ The neighbor set of a node in the information graph.
+$\mathcal{N}(\cdot)$ The neighbor set of a node in the weighted graph.
 $p_{extract}$ The extraction instruction prompt.
@@
 $\mathcal{E}_{oi}^{t}$ Edges added based on text $t$.
-Figure 2.
-An overview of the GORAG model is as follows.
-In Part 1, GORAG constructs a graph based on keywords extracted from the few-shot training data, taking into account the keyword extraction confidence and relatedness to the target labels.
-In Part 2, GORAG performs adaptive graph retrieval and utilizes the retrieved graph information to filter the target labels.
-In Part 3, GORAG carry out classification with the testing texts and the candidate labels retrieved from the information graph.
-In Part 4, GORAG applies online indexing to update the graph with newly extracted nodes from testing texts through online indexing.
-2.2.3. LLM-based models
-Recently, Large Language Model (LLM)-based models have undergone rapid development (Fernandez et al., 2023; Amer-Yahia et al., 2023; Zhang et al., 2024b; Miao et al., 2024) and have been successfully adapted to various tasks, including data discovery (Arora et al., 2023; Dong et al., 2023; Kayali et al., 2024), entity or schema matching (Zhang et al., 2023b; Tu et al., 2023; Fan et al., 2024; Du et al., 2024), and natural language to SQL conversion (Li et al., 2024; Ren et al., 2024; Trummer, 2022; Gu et al., 2023).
-Notably, LLMs are inherently capable of inference without fine-tuning (Sun et al., 2021; Jiang et al., 2023; Touvron et al., 2023; Achiam et al., 2023; Yang et al., 2024), making them originally suitable for dynamic text classification tasks.
-However, the lack of fine-tuning can lead LLMs to generate incorrect answers, as they lack task-specific knowledge, these incorrect answers are often referred to as hallucinations (Zhang et al., 2023c).
-To mitigate hallucinations, researchers have provided LLMs with side information for classification, namely long context RAG models.
-For example, Propositionizer (Chen et al., 2023) applies a fine-tuned LLM to convert side information into atomic expressions, to facilitate fine-grained information retrieval.
-RAPTOR (Sarthi et al., 2024) clusters side information and then generate summaries for each cluster, to help with the model prediction.
-However, the retrieved contents from these models remain unstructured and can be lengthy, which impedes the efficiency and effectiveness of LLMs, leading to lost-in-the-middle issue (Liu et al., 2024).
-To address this, compression-based RAG models were proposed, these models try to compress side information to reduce the input length.
-Based on how they compress model inputs, these models can be classified into prompt compressor models, and graph-based RAG models.
-On the one hand, prompt compressor models, such as LLMLingua2 (Pan et al., 2024), apply LLM’s generation perplexity to filter out un-important tokens in the model input.
-Based on LLMLingua2, LongLLMLingua (Jiang et al., 2024) further considers the instruction to the model when compression, make the model instruction-aware.
-On the other hand, graph-based RAG models index side informations into a graph, and retrieve graph components or summaries, which are shorter than long contexts thet traditional RAG models retrieved.
-For instance, GraphRAG (Edge et al., 2024) constructs a graph and aggregates nodes into graph communities, then generate community summaries to help the LLM prediction.
-However, GraphRAG requires re-creating the community after each graph index, which incurs significant computational overhead.
-To address this limitation of GraphRAG, two alternative models, HippoRAG (Gutiérrez et al., 2024) and LightRAG (Guo et al., 2024), have been proposed, which skip the formulation of communities and directly retrieve nodes or pathes from graphs.
-HippoRAG (Gutiérrez et al., 2024) retrieve nodes whose retrieval score is above a human defined threshold.
-However, the optimal retrieval threshold can vary across different data samples, and a globally fixed threshold selected by human may be suboptimal for the entire dataset.
-To address this issue, LightRAG (Guo et al., 2024) calculates the embedding similarity of text extracted entities with graph nodes, to achieve a one-to-one mapping and retrieve all triples involved these graph nodes.
-3. Methdology
-To address the aforementioned issues, we propose GORAG, a novel approach that achieves adaptive retrieval by extracting valuable side information from a minimum-cost spanning tree generated on the constructed graph.
-Specifically, GORAG takes into account each keyword’s with in the training and texting text corpus and relatedness to the class label by assigning different costs to edges on the constructed graph, allowing for more informed retrieval decisions.
-Moreover, GORAG incorporates an online indexing mechanism, which enables the model to index valuable information extracted from the testing texts to the constructed graph in real-time, thereby enhancing its ability to make accurate predictions in the future.
-3.1. Framework Overview
-As illustrated in Figure 2, our model consists of four primary components, i.e., Graph Construction, Graph Retrieval, Classification, and Online Indexing.
-Figure 3.
-An example of the constructed information graph from WOS dataset, where red nodes denote the label node and black nodes denote the keyword node.
-In this subgraph, node Entity and node Neighborhood are considered most related to the label Relational Database, the edge between them with the label have the lowest cost assigned by our edge weighting mechanism.
-Part 1: Graph Construction Graph construction proposes is to construct or update the information graph $\mathcal{G}^{r}_{new}(\mathcal{V}^{r}_{new},\mathcal{E}^{r}_{new},\mathcal{W}^{r}_{new})$ at round $r$ based on the labeled data $\mathcal{D}_{new}^{r}=\cup_{y_{i}\in\mathcal{Y}_{\text{new}}^{r}}\{t_{j},y_{i}\}_{j=1}^{k}$ for at the $r$-th round.
-A small subgraph of the information graph created on WOS dataset is shown in Figure 3, graph formed like this will be used to provide retrieved-augmented information as context for unseen text, enabling LLMs to better understand the unseen text and accurately predict its labels.
-Specifically, under each new data $\mathcal{D}_{new}^{r}$ at the $r$-th round, we first extract the new keywords $\mathcal{K}_{new}^{r}\in\{t_{j}\}_{j=1}^{|\mathcal{D}_{new}^{r}|}$ from the few-shot training texts $\mathcal{T}^{r}_{new}=\{t_{j}\}_{j=1}^{|\mathcal{D}_{new}^{r}|}$.
-Then, we will then assign an edge $e^{r}_{v,y}\in\mathcal{E}^{r}_{new}$ between each keyword $v\in\mathcal{K}_{new}^{r}$ and each label $y\in\mathcal{Y}^{r}_{new}$.
-We then compute the weight $w^{r}_{v,y}\in\mathcal{W}^{r}_{new}$ for each edge $e^{r}_{v,y}\in\mathcal{E}^{r}_{new}$ based on the keyword’s importance with in the text corpus and relatedness to the label $y$.
-We then can merge all graphs $\mathcal{G}^{r}_{new}(\mathcal{V}^{r}_{new},\mathcal{E}^{r}_{new},\mathcal{W}^{r}_{new})$ at each each round $r$ as the full graph $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r},\mathcal{W}^{r})$.
-More details can refer to Section 3.2.
-Part 2: Graph Retrieval The Graph Retrieval process maps the extracted keyword nodes $\mathcal{V}^{t}$ from the unlabeled text $t$ to the constructed information graph $\mathcal{G}^{r}$ and generates a minimum-cost spanning tree on $\mathcal{G}^{r}$ that includes all these keywords.
-From this minimum-cost spanning tree, the candidate label set $\hat{\mathcal{Y}}_{t}^{r}$ is obtained, which is a reduced subset of the original target label set $\mathcal{Y}^{r}$.
-For more details, please refer to Section 3.3.
-Part 3: Classification After retrieving the graph and obtaining the candidate types $\hat{\mathcal{Y}}_{t}^{r}$ for the text $t$ at round $r$, the classification procedure involves performing the final classification based on a large language model.
-This process utilizes the text $t$, the candidate labels $\hat{\mathcal{Y}}_{t}^{r}$ retrieved from the information graph $\mathcal{G}^{r}$, and descriptions $\mathcal{K}_{y_{i}}$ associated with each candidate label $y_{i}\in\hat{\mathcal{Y}}_{t}^{r}$.
-For more details, please refer to Section 3.4.1.
-Part 4: Online Indexing The online indexing procedure dynamically indexes keywords, denoted as $\mathcal{V}_{\text{notexist}}^{t}$, which extracted from the testing text $t$ but are not in the existing graph $\mathcal{G}^{r}$.
-We then integrate these keywords into the graph to further enrich its structure, with weights assigned based on their importance within the text corpus and their relatedness to the predicted label $y^{*}_{t}$ for the testing text $t$.
-This process enhances the model’s ability to make more accurate predictions in the future.
-For more details, please refer to Section 3.4.2.
-3.2. Part 1: Graph Construction
-In this subsection, we would introduce the graph construction procedure of GORAG.
-To generate the information graph $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r},\mathcal{W}^{r})$ for each the $r$-th round, GORAG applies multiple instructions to the LLM for creating the graph node set $\mathcal{V}^{r}$ and the edges $e^{r}_{v,y}\in\mathcal{E}^{r}$ between keyword node $v$ and label node $y$.
-Also, we will assign weight $w^{r}_{v,y}\in\mathcal{W}^{r}$ for each each edge $e(v,y)$.
-The pseudo code of the graph construction algorithm is shown in Algorithm 1.
-Algorithm 1 Graph Construction Algorithm of GORAG at the $r$-th round
-0:  $\mathcal{D}_{new}^{r}$: Training text set at the $r$-th round and the provided label $y_{i}\in\mathcal{Y}_{new}^{r}$ for each training text $t_{j}$; $p_{extraction},\ p_{gen}$: The extraction and description generation instruction prompt; $\mathcal{G}^{r-1}$: Information graph of the previous round $r-1$.
-0:  $\mathcal{G}^{r}$: Constructed information graph at the $r$-th round.
+Appendix B Complexity Analysis
+In this section, we analysis the time and space complexity of GORAG’s graph construction, retrieval and prediction procedure.
+We denote the maximum number of terms of the input text as $m_{t}$, the number of unique terms of training corpus $\mathcal{D}_{new}^{r}$ as $u$, the LLM’s maximum input token length as $m_{i}$, LLM’s its maximum extraction, generation, and classification token length as $m_{e}$, $m_{g}$, and $m_{c}$, respectively.
+B.1. Graph Construction Complexity
+Firstly, we analysis the time and space complexity of GORAG’s graph construction.
+For the text keyword extraction procedure, the time complexity would be $O(|\mathcal{D}_{new}^{r}|(m_{i}+m_{e}))$; If label names are available, the time complexity of generat label descriptions and extract label keywords would be $O(|\mathcal{Y}^{r}_{new}|(m_{i}+m_{g}+m_{e}))$; Calculating the TFIDF and indexing edges to graph would require $O(|\mathcal{V}_{new}^{r}|(|\mathcal{D}_{new}^{r}|m_{t}+u))$ times, and merging graph requires
+$O(|\mathcal{Y}^{r}_{new}||\mathcal{Y}^{r-1}|)$. Hence, the total time complexity of GORAG’s graph construction at the $r$-th round would be:
+$O(|\mathcal{D}_{new}^{r}|(m_{i}+m_{e})+|\mathcal{Y}^{r}_{new}|(m_{i}+m_{g}+m_{e}+|\mathcal{Y}^{r-1}|)+|\mathcal{V}_{new}^{r}|(|\mathcal{D}_{new}^{r}|m_{t}+u))$.
+For the space complexity, the graph is stored with the weighted adjency matrix, hence needs $O(|\mathcal{E}^{r}|)$ space; Storing the training corpus at the $r$-th round would need $O(u|\mathcal{D}_{new}^{r}|)$ space; Storing the representive keywords $\mathcal{K}_{\hat{\mathcal{Y}_{t}^{r}}}$ would need $O(|\mathcal{K}_{\hat{\mathcal{Y}_{t}^{r}}}|)$ space.
+Hence the total space complexity of GORAG’s graph construction at the $r$-th round would be:
+$O(|\mathcal{E}^{r}|+u|\mathcal{D}_{new}^{r}|+|\mathcal{K}_{\hat{\mathcal{Y}_{t}^{r}}}|)$.
+B.2. Retrieval and Classification Complexity
+The time complexity of GORAG’s adaptive candidate type generation algorithm is the same with the Mehlhorn algorithm, which is $O(|\mathcal{E}^{r}|+|\mathcal{V}^{r}|log|\mathcal{V}^{r}|)$ (Mehlhorn, 1988b); The time complexity of the online indexing mechanism would cost $|\mathcal{V}_{notexist}^{r}|(|\mathcal{T}_{test}^{r}|m_{t}+u)$; The time complexity of the final classification by LLM is $O(m_{i}+m_{c})$.
+Hence, the total time complexity of GORAG’s adaptive retrieval and classification can be denoted as
+$O(|\mathcal{T}_{test}^{r}|(|\mathcal{E}^{r}|+|\mathcal{V}_{notexist}^{r}|(|\mathcal{T}_{test}^{r}|m_{t}+u)+|\mathcal{V}^{r}|log|\mathcal{V}^{r}|+m_{i}+m_{c})).$
+For the space complexity, $O(|\mathcal{E}^{r}|)$ to store the graph, and $O(u|\mathcal{T}_{test}^{r}|)$ space is needed to store the testing corpus.
+Hence, the total space complexity of GORAG’s Retrieval and Classification procedure is
+$O(|\mathcal{E}^{r}|+u|\mathcal{T}_{test}^{r}|).$
+Appendix C Pseudo Code of GORAG
+In this section, we provide the pseudo code of GORAG’s graph construction and adaptive candidate type generation algorithm in 1 and 2, respectively.
+Appendix D Hardness Proof of The Adaptive Candidate Type Generation problem
+Theorem D.1.
+The Adaptive Candidate Type Generation problem is NP-hard.
+Proof.
+To demonstrate that the Adaptive Candidate Type Generation problem is NP-hard, we provide a simple reduction of our problem from the Steiner Tree problem.
+Since the Steiner Tree problem is already proven to be NP-hard (Su et al., 2020), we show that there is a solution for the Steiner Tree problem if and only if there is a solution for our problem.
+Firstly, given a solution $S$ for our problem, we can construct a Steiner Tree by generating a minimum spanning tree for all nodes in $S$ on graph $\mathcal{G}^{r}$ then connecting all nodes from $\mathcal{V}_{exist}^{t}$ to their closest neighbor nodes in $S$.
+Secondly, any Steiner Tree $ST$ that contains any node $y\in\mathcal{V}^{r}_{t}$ is also a solution of our problem with |$|\{y\in\mathcal{V}^{r}_{t}\}|$ nodes.
+Thus, we prove that the adaptive candidate type generation problem is NP-hard.
+∎
+Algorithm 1 Graph Construction Algorithm of GORAG at round $r$
+0:  $\mathcal{D}_{new}^{r}$: Training text set at the $r$-th round and the provided label $y_{i}\in\mathcal{Y}_{new}^{r}$ for each training text $t_{j}$; $p_{extraction},\ p_{gen}$: The extraction and description generation instruction prompt; $\mathcal{G}^{r-1}$: Weighted graph of the previous round $r-1$.
+0:  $\mathcal{G}^{r}$: Constructed weighted graph at the $r$-th round.
 1:  Let $\mathcal{V}_{text}^{r}=\emptyset$, $\mathcal{V}_{label}^{r}=\emptyset$, $\mathcal{E}_{new}^{r}=\emptyset$
@@
 18:  return  $\mathcal{G}^{r}$
-At each $r$-th round, given the labeled training text $\mathcal{D}_{new}^{r}=\cup_{y_{i}\in\mathcal{Y}_{\text{new}}^{r}}\{t_{j},y_{i}\}_{j=1}^{k}$, GORAG first extracts text keywords $\mathcal{K}^{r}_{new}$ from the text $\mathcal{T}^{r}_{new}=\cup_{y_{i}\in\mathcal{Y}_{\text{new}}^{r}}\{t_{j}\}_{j=1}^{k}$ to serve as graph nodes.
-Specifically, we use the LLM model with an extraction instruction prompt $p_{extract}$, such as $p_{extract}=$Please extract some keywords from the following passage”.
-Then, we can get all text keywords $\mathcal{K}^{r}_{new}$ based on texts $\mathcal{T}^{r}_{new}$ as follows.
-(2) $\mathcal{K}^{r}_{new}=\bigcup_{t\in\mathcal{T}^{r}_{new}}LLM(p_{extract},t).$
-Also, by incorporating the new candidate labels $\mathcal{Y}^{r}_{new}$ at the $r$-th round, the graph node set $\mathcal{V}^{r}_{new}$ at the $r$-th round can be obtained as follows.
-(3) $\mathcal{V}_{new}^{r}=\mathcal{K}_{new}^{r}\cup\mathcal{Y}_{new}^{r}.$
-Then, GORAG links each keyword node $v\in\mathcal{K}_{{new}}^{r}$ to its corresponding label node $y\in\mathcal{Y}_{{new}}^{r}$, indicating that the keyword $v$ appears in texts associated with the label $y$.
-We discuss how to compute the weight between each keyword node $v\in\mathcal{K}_{{new}}^{r}$ and each label $y\in\mathcal{Y}_{{new}}^{r}$.
-Considering the keywords $\mathcal{K}_{new}^{r}$ from the labeled text are not uniformly related to the text’s label $\mathcal{Y}_{new}^{r}$, we apply an edge weighting mechanism to assign a weight $w_{v,y}$ to each keyword-label link.
-This weight can be regarded as the correlation between keyword and each label.
-Firstly, we apply the $l_{2}$ normalized TF-IDF score (Sammut and Webb, 2010) to measure the importance and relatedness of a particular keyword $v\in\mathcal{K}_{new}^{r}$ and label $y\in\mathcal{Y}_{new}^{r}$ of the text $t\in\mathcal{T}^{r}$, where $\mathcal{T}^{r}=\cup_{i=1}^{r}\mathcal{T}_{new}^{i}$ denotes texts seen so far.
-Formally, given a keyword node $v\in\mathcal{K}_{new}^{r}$ and one text $t\in\mathcal{T}^{r}$, the correlation score $CS(v,t)$ between $v$ and $t$ is defined as follows.
-(4) $CS(v,t)=\frac{count(v,t)}{|t|}\times log\frac{|\mathcal{T}^{r}|}{1+|t_{j}:v\in t_{j},\ t_{j}\in\mathcal{T}^{r}|},$
-where $count(v,t)$ is the number of times that the term $v$ appears in the text $t$, and $|t_{j}:v\in t_{j},\ t_{j}\in\mathcal{T}^{r}|$ denotes the number of texts in the corpus $\mathcal{T}^{r}$ that contain the keyword $v$.
-As the keyword $v$ can be extracted from multiple text source and from different rounds, the final edge weight $w_{v,y}^{r}$ of edge $e_{v,y}$ at the $r$-th round is calculated an the average of weights from all these texts.
-Formally, given the keyword $v\in\mathcal{K}_{new}^{r}$, label $y\in\mathcal{Y}_{new}^{r}$, and the texts $\mathcal{T}^{r,y}=\{t_{i}\}_{i=1}^{|\mathcal{T}^{r,y}|}$ consisting of texts $t_{i}$ associated with label $y$, the weight $w_{v,y}^{r}$ between $v$ and $y$ at the $r$-th round can be computed as follows:
-(5) $w_{v,y}^{r}=\frac{\sum_{t_{j}\in\mathcal{T}^{r,y,v}}1-CS(v,t_{j})}{|\mathcal{T}^{r,y,v}|},$
-where $\mathcal{T}^{r,y,v}=\{t_{j}|v\in t_{j}\land t_{j}\in\mathcal{T}^{r,y}\}$ is the text associated with label $y$ and containing the keyword $v$.
-We denote the generated graph for $\mathcal{K}_{new}^{r}$ and $\mathcal{Y}_{new}^{r}$ at the $r$-th round as $\mathcal{G}^{r}_{new}(\mathcal{V}_{new}^{r},\mathcal{E}_{new}^{r},\mathcal{W}_{new}^{r})$, where the node $\mathcal{V}^{r}_{new}=\mathcal{K}_{new}^{r}\cup\mathcal{Y}_{new}^{r}$ and edge $e^{r}_{v,y}\in\mathcal{E}_{new}^{r}$ is associated with the weight $w^{r}_{v,y}\in\mathcal{W}_{new}^{r}$.
-We then discuss how to merge $\mathcal{G}^{r}_{new}$ into the graph from previous rounds $\mathcal{G}^{r-1}(\mathcal{V}^{r-1},\mathcal{E}^{r-1})$ to form the full graph $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r})$ as follows.
-(6) $\mathcal{G}^{r}=Merge(\mathcal{G}^{r}_{new},\mathcal{G}^{r-1}),$
-where $\mathcal{V}^{r}=\mathcal{V}^{r}_{new}\cup\mathcal{V}^{r-1}$, $\mathcal{E}^{r}=\mathcal{E}^{r}_{new}\cup\mathcal{E}^{r-1}$, and $\mathcal{W}^{r}=\mathcal{W}^{r}_{new}\cup\mathcal{W}^{r-1}$.
-Particularly, $\mathcal{G}^{0}=\emptyset$ and $\ r\geq 1$, Also, to guarantee the graph connectivity of the resulting graph, we add new edges between every newly added label node $y_{n}\in\mathcal{Y}_{new}^{r}$ and each old label node $y_{o}\in\mathcal{Y}^{r-1}$, the edge weight $w_{v,y}^{r}$ of edge $e_{y_{n},y_{o}}$ at round $r$ is calculated with the average weight of all edges that link keywords with label node $y_{n}$ or $y_{o}$, respectively:
-(7) $w_{y_{n},y_{o}}^{r}=\frac{AvgWei(y_{n},\mathcal{G}^{r})+AvgWei(y_{o},\mathcal{G}^{r-1})}{2}$
-(8) $AvgWei(y,\mathcal{G}^{r})=\frac{\sum_{v\in\mathcal{M}_{y}^{r}}w_{v,y}^{r}}{2\times|\mathcal{M}_{y}^{r}|},$
-where $\mathcal{N}^{r}(y)$ denote the neighbor node set of label node $y$ in graph $\mathcal{G}^{r}$.
-Also, $\mathcal{M}_{y}^{r}=\{v\ |\ v\in\mathcal{N}^{r}(y)\ \land\ v\notin\mathcal{Y}^{r}\}$ denotes all neighbor nodes of label node $y$ that represent keywords, but not labels at the $r$-th round.
-After merge, the graph $\mathcal{G}^{r}$ would be used for future retrieval, and be further updated by GORAG’s online indexing mechanism later in Section 3.3.
+Appendix E Experiment Details
+E.1. Further introduction of chosen baselines
+NN-based Models
+  • Entailment (Xia et al., 2021): Entailment concatenates the text data with each of the label names to form multiple entailment pairs with one text sample, hence increasing the number of training data and enhance its finetuning of a RoBERTa PLM (Liu, 2019).
+Entailment then carry out classification based on these entailment pairs and convert the text classification task into a binary classification task, for detecting whether the entailment pair formed is correct.
+When all entailment pairs of a particular text are considered to be incorrect, Entailment will directly give up the classification of this text.
+In this paper, we consider this classification as a hallucination.
+Long Context RAG Models
+  • NaiveRAG (Gao et al., 2023): NaiveRAG acts as a foundational baseline of current RAG models.
+When indexing, it stores text segments of the labeled texts in a vector database using text embeddings.
+When querying, NaiveRAG generates query texts’ vectorized representations to retrieve side information based on the highest similarity in their embeddings.
+  • Propositionizer (Chen et al., 2023): Propositionizer applies a fine-tuned LLM to convert side information into atomic expressions, namely propositions, to facilitate more fine-grained information retrieval than NaiveRAG.
+Compression RAG Models
+  • LongLLMLingua (Jiang et al., 2024a): LongLLMLingua is a instruction aware prompt compressor model, it applies LLM’s generation perplexity to filter out un-important tokens of the model input based on the retrieved side information and the task instruction.
+  • GraphRAG (Edge et al., 2024): GraphRAG is a graph-Based RAG model that employs the LLM to extract texts’ entities and relations, which are then represented as nodes and edges in the weighted graph.
+GraphRAG then aggregates nodes into communities, and generates a comprehensive community report to encapsulate global information from texts.
+  • LightRAG (Guo et al., 2024): LightRAG skips the GraphRAG’s formulation of graph communities and directly retrieve nodes or pathes from created graphs.
+It calculates the embedding similarity of text extracted entities with graph nodes, to achieve a one-to-one mapping from keywords to graph nodes and retrieve all triples involved these nodes.
 Algorithm 2 Adaptive Candidate Type Generation Algorithm
-0:  $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r},\mathcal{W}^{r})$: The constructed information graph; $\mathcal{V}_{exist}^{t}\subseteq\mathcal{V}^{r}$: A set of keyword nodes extracted from text $t$ and can be mapped to graph $\mathcal{G}$; $\mathcal{Y}^{r}$: The target label set at the $r$-th round.
+0:  $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r},\mathcal{W}^{r})$: The constructed weighted graph; $\mathcal{V}_{exist}^{t}\subseteq\mathcal{V}^{r}$: A set of keyword nodes extracted from text $t$ and can be mapped to graph $\mathcal{G}$; $\mathcal{Y}^{r}$: The target label set at the $r$-th round.
 0:  $\hat{\mathcal{Y}}^{r}_{t}$: The candidate type retrieved for text $t$.
@@
 14:  return  $\hat{\mathcal{Y}}^{r}_{t}$
-3.3. Part 2: Graph Retrieval
-In this subsection, we would introduce the graph retrieval procedure of GORAG.
-With the graph constructed in Part 1, GORAG is able to carry out adaptive reteieval algorithm with the keywords extracted from the testing texts.
-To begin with, GORAG extract keywords $\mathcal{V}_{test}^{t}$ for each testing text $t\in T_{test}$ in the same manner with Equation (2), then, $\mathcal{V}_{test}^{t}$ would be splitted into two subsets:
-(9) $\mathcal{V}_{exist}^{t}\cup\mathcal{V}_{notexist}^{t}=\mathcal{V}_{test}^{t},\ \mathcal{V}_{exist}^{t}\cap\mathcal{V}_{notexist}^{t}=\emptyset,$
-where $\mathcal{V}_{exist}^{t}$ and $\mathcal{V}_{notexist}^{t}$ denotes the keywords in $\mathcal{V}_{test}^{t}$ that already exist and not yet exist in $\mathcal{G}^{r}$ at the current the $r$-th round respectively.
-Later, $\mathcal{V}_{exist}^{t}$ would be applied for achieving adaptive retrieval, and $\mathcal{V}_{notexist}^{t}$ would be applied for online indexing to further enrich the constructed graph (Further illustrated in Section 3.4).
-To achieve the adaptive retrieval, GORAG try to find the minimum cost spanning tree that contain all keyword nodes within $\mathcal{V}_{exist}^{t}$.
-The intuition behind this approach is that a minimum cost spanning tree (MST) spans the entire graph to cover all given nodes with the smallest possible spanning cost.
-Consequently, nodes within the generated MST can be considered important for demonstrating the features of the given node set.
-The generation of an MST is a classical combinatorial optimization problem with an optimal solution determined solely by the set of given nodes.
-By generating the MST, we eliminate the need for any human-defined thresholds, relying instead on the LLM-extracted keywords to form the spanning tree.
-Definition 3.1 (Adaptive Candidate Type Generation Problem).
-Given an undirected weighted information graph $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r},\mathcal{W}^{r})$, a set of keywords $\mathcal{V}_{exist}^{t}$ extracted from text $t$ that can be mapped to nodes in $\mathcal{V}^{r}$, and the target label set $\mathcal{Y}^{r}$ at the $r$-th round, our target is to find a set of labels nodes ${\mathcal{Y}}^{r}_{t}\in{\mathcal{Y}}^{r}$.
-Firstly, we identify a subgraph $\mathcal{G}^{r}_{t}(\mathcal{V}^{r}_{t},\mathcal{E}^{r}_{t},\mathcal{W}^{r}_{t})$ of $\mathcal{G}^{r}$ by minimizing the edge weight sum as follows.
-(10) $\min_{e^{r}_{u.v}\in\mathcal{E}^{r}_{t}}w^{r}_{u,v}$
-(11) $s.t.$ $v\in\mathcal{V}^{r}_{t},\forall v\in\mathcal{V}_{exist}^{t}$
-Then, since the subgraph nodes $\mathcal{V}^{r}_{t}$ contains both keyword nodes and labels, we take the label nodes ${\mathcal{Y}}^{r}_{t}\in\mathcal{V}^{r}_{t}$ as our target candidate nodes for the text $t$.
-Theorem 3.2.
-The Adaptive Candidate Type Generation problem is NP-hard.
-Proof.
-To demonstrate that the Adaptive Candidate Type Generation problem is NP-hard, we provide a simple reduction of our problem from the Steiner Tree problem.
-Since the Steiner Tree problem is already proven to be NP-hard (Su et al., 2020), we show that there is a solution for the Steiner Tree problem if and only if there is a solution for our problem.
-Firstly, given a solution $S$ for our problem, we can construct a Steiner Tree by generating a minimum spanning tree for all nodes in $S$ on graph $\mathcal{G}^{r}$ then connecting all nodes from $\mathcal{V}_{exist}^{t}$ to their closest neighbor nodes in $S$.
-Secondly, any Steiner Tree $ST$ that contains any node $y\in\mathcal{V}^{r}_{t}$ is also a solution of our problem with |$|\{y\in\mathcal{V}^{r}_{t}\}|$ nodes.
-Thus, we prove that the adaptive candidate type generation problem is NP-hard.
-∎
-Since this problem is the NP-hard, it is infeasible to obtain the optimal result in polynomial time.
-Therefore, to solve this problem, we propose a greedy algorithm, which generates the candidate type set ${\mathcal{Y}}^{r}_{t}\subseteq\mathcal{Y}^{r}$ for text $t$ at the $r$-th round.
-The detail of our algorithm is shown in Algorithm 2.
-Firstly, we calculate the shortest path between each keyword node $v$ in $\mathcal{V}_{exist}^{t}$ to all other nodes in $\mathcal{G}^{r}$ by calculating the minimum spanning tree $MST$ of $\mathcal{G}^{r}$ (line 1-2).
-Here, the keyword nodes $v$ in $\mathcal{V}_{exist}^{t}$ are served as terminal nodes that determins the final genrated candidate types w.r.t.
-to the information graph $\mathcal{G}^{r}$ Secondly, we create a new auxiliary graph $H$ where the edges represent the shortest paths between the closest terminal nodes (line 3-7).
-Thirdly, we construct the minimum spanning tree $MST^{\prime}$ of the auxiliary graph $H$ (line 9), and then we add the shortest paths between each two nodes in $MST^{\prime}$ to the Steiner Tree $ST$ (line 10-12).
-Lastly our candidate types $\hat{\mathcal{Y}}^{r}_{t}$ for text $t$ are calculated as the interselect of $ST$ and all target labels $\mathcal{Y}^{r}$ at round $r$: $ST\cap\mathcal{Y}^{r}$ (line 13).
-Note that in GORAG’s retrieval algorithm, the minimum-cost spanning tree is generated solely based on the constructed graph and the keywords extracted from the text.
-This approach eliminates the need for any manually defined retrieval thresholds, enabling GORAG to perform fully adaptive and context-driven retrieval.
-3.4. Part 3: Classification and Online Indexing
-In this subsection, we introduce GORAG’s two main components, i.e., classification and online indexing.
-For classification, GORAG utilizes an LLM to predict the class label for each unseen test text by retrieving adaptive demonstrations.
-For online indexing, GORAG dynamically updates the graph $\mathcal{G}^{r}$ by incorporating unseen keywords from the test text $t$ as new nodes and connecting them to the predicted label $y^{*}_{t}$ with weighted edges.
-This iterative process enriches the graph and enhances the framework’s ability to make accurate predictions in the future.
-3.4.1. Classification
-In this part, we introduce how GORAG performs text classification based a large language model.
-Specifically, for each unlabeled text $t$, GORAG predicts its class label $y^{*}$ by constructing an input $c_{t}$ for the LLM, which captures contextual and structural information about the text and its candidate labels.
-Specifically, given each unlableed text $t$, we will use LLM to predict its class label $y^{*}$.
-We first generate the LLM input $c_{t}$ as follows.
-(12) $c_{t}=Concat(t,\mathcal{K}^{t},\hat{\mathcal{Y}}^{r}_{t},\mathcal{K}_{\hat{\mathcal{Y}}^{r}_{t}}),$
-which is the text concatenation of the extracted keywords $\mathcal{K}^{t}$ from the text $t$, the candidate labels $\hat{\mathcal{Y}}^{r}_{t}$ obtained by Algorithm 2.
-Also, $\mathcal{K}_{\hat{\mathcal{Y}}^{r}_{t}}=\{\mathcal{K}_{y_{i}}\}_{i=1}^{|\hat{\mathcal{Y}}^{r}_{t}|}$ and $\mathcal{K}_{y_{i}}$ is the representative keywords of each label $y_{i}\in\hat{\mathcal{Y}}^{r}_{t}$.
-(13) $\mathcal{K}_{y_{i}}=LLM(p_{gen},\mathcal{D}^{r}),$
-and a classification instruction prompt $p_{classify}$.
-(14) $y^{*}_{t}=LLM(p_{classify},c_{t}),$
-As a result, the LLM would try to select the best-suited label $y^{*}_{t}\in\mathcal{Y}^{t}$ to annotate the text $t$.
-3.4.2. Online indexing
-To fully leverage the text-extracted keywords, GORAG utilizes an online indexing mechanism to incrementally update keywords that do not yet exist in the information graph $\mathcal{G}^{r}$ at the $r$-th round to the information graph $\mathcal{G}^{r}$ based on the text $t$’s predicted label $y^{*}_{t}$.
-To be specific, each keyword node $v\in\mathcal{V}_{notexist}^{t}$ would be directed added to the original graph’s node set $\mathcal{V}^{r}$ and be assigned with an edge $e_{v,y}$ connect it with the predicted label $y$:
-(15) $\mathcal{V}^{r}=\mathcal{V}^{r}\cup\mathcal{V}_{notexist}^{t},\ \mathcal{E}^{r}=\mathcal{E}^{r}\cup\mathcal{E}_{oi}^{t},$
-where $\mathcal{E}_{oi}^{t}=\{e_{v,y_{t}}|v\in\mathcal{V}_{notexist}^{t}\}$ denotes the set of all newly assigned edges $e_{v,y}$ between keyword node $v$ and its predicted label $y_{t}^{*}$, for these newly assigned edges, their weight is calculated with the edge weighting mechanism illustrated in Equation (5).
-Table 2. Statistics of the WOS and Reuters dataset.
-Dataset Avg. Text Token R1 R2 R3 R4 Total
-Testing data Label # Testing data Label # Testing data Label # Testing data Label # Testing data
-WOS (Kowsari et al., 2017) 200 2,417 32 2,842 53 2,251 30 1,886 18 9,396
-Reuters (Bao et al., 2020) 168 80 8 80 8 80 8 70 7 310
-This online indexing mechanism serves two main purposes.
-First, it incrementally enriches the information graph $\mathcal{G}^{r}$ by incorporating new, previously unseen keywords, thus expanding the graph’s vocabulary and improving its adaptability.
-Second, by linking these new keywords to their predicted labels with weighted edges, the mechanism captures their relevance and context more effectively, enabling better utilization of the graph for future predictions.
-3.5. Complexity Analysis
-In this subsection, we analysis the time and space complexity of GORAG’s graph construction, retrieval and prediction procedure.
-We denote the maximum number of terms of the input text as $m_{t}$, the number of unique terms of training corpus $\mathcal{D}_{new}^{r}$ as $u$, the LLM’s maximum input token length as $m_{i}$, LLM’s its maximum extraction, generation, and classification token length as $m_{e}$, $m_{g}$, and $m_{c}$, respectively.
-Graph Construction Complexity Firstly, we analysis the time and space complexity of GORAG’s graph construction.
-For the text keyword extraction procedure, the time complexity would be $O(|\mathcal{D}_{new}^{r}|(m_{i}+m_{e}))$; If label names are available, the time complexity of generat label descriptions and extract label keywords would be $O(|\mathcal{Y}^{r}_{new}|(m_{i}+m_{g}+m_{e}))$; Calculating the TFIDF and indexing edges to graph would require $O(|\mathcal{V}_{new}^{r}|(|\mathcal{D}_{new}^{r}|m_{t}+u))$ times, and merging graph requires $O(|\mathcal{Y}^{r}_{new}||\mathcal{Y}^{r-1}|)$.
-Hence, the total time complexity of GORAG’s graph construction at the $r$-th round would be:
-$O(|\mathcal{D}_{new}^{r}|(m_{i}+m_{e})+|\mathcal{Y}^{r}_{new}|(m_{i}+m_{g}+m_{e}+|\mathcal{Y}^{r-1}|)+|\mathcal{V}_{new}^{r}|(|\mathcal{D}_{new}^{r}|m_{t}+u))$.
-For the space complexity, the graph is stored with the weighted adjency matrix, hence needs $O(|\mathcal{E}^{r}|)$ space; Storing the training corpus at the $r$-th round would need $O(u|\mathcal{D}_{new}^{r}|)$ space; Storing the representive keywords $\mathcal{K}_{\hat{\mathcal{Y}_{t}^{r}}}$ would need $O(|\mathcal{K}_{\hat{\mathcal{Y}_{t}^{r}}}|)$ space.
-Hence the total space complexity of GORAG’s graph construction at the $r$-th round would be:
-$O(|\mathcal{E}^{r}|+u|\mathcal{D}_{new}^{r}|+|\mathcal{K}_{\hat{\mathcal{Y}_{t}^{r}}}|)$.
-Retrieval and Classification Complexity The time complexity of GORAG’s adaptive candidate type generation algorithm is the same with the Mehlhorn algorithm, which is $O(|\mathcal{E}^{r}|+|\mathcal{V}^{r}|log|\mathcal{V}^{r}|)$ (Mehlhorn, 1988); The time complexity of the online indexing mechanism would cost $|\mathcal{V}_{notexist}^{r}|(|\mathcal{T}_{test}^{r}|m_{t}+u)$; The time complexity of the final classification by LLM is $O(m_{i}+m_{c})$.
-Hence, the total time complexity of GORAG’s adaptive retrieval and classification can be denoted as
-$O(|\mathcal{T}_{test}^{r}|(|\mathcal{E}^{r}|+|\mathcal{V}_{notexist}^{r}|(|\mathcal{T}_{test}^{r}|m_{t}+u)+|\mathcal{V}^{r}|log|\mathcal{V}^{r}|+m_{i}+m_{c})).$
-For the space complexity, $O(|\mathcal{E}^{r}|)$ to store the graph, and $O(u|\mathcal{T}_{test}^{r}|)$ space is needed to store the testing corpus.
-Hence, the total space complexity of GORAG’s Retrieval and Classification procedure is
-$O(|\mathcal{E}^{r}|+u|\mathcal{T}_{test}^{r}|).$
-4. Experiments
-In this section, we present the experimental evaluation of our framework GORAG on the DFSTC task.
-We compare the performance of GORAG against six effective baselines spanning three technical categories, using two datasets with distinct characteristics.
-Specifically, we first outline the experimental setup in Section 4.1, including details on datasets, baselines, evaluation metrics, and hyperparameter configurations.
-Next, we report the experimental results, focusing on both effectiveness and efficiency evaluations, in Section 4.2.
-Finally, we conduct an ablation study and a case study, presented in Section 4.3 and Section 4.4, respectively.
-Table 3. Experiment results on WOS and Reuters dataset.
-Dataset Category Model Round
-R1 R2 R3 R4
-1-shot 5-shot 10-shot 1-shot 5-shot 10-shot 1-shot 5-shot 10-shot 1-shot 5-shot 10-shot
-WOS NN-based Entailment 0.3695 0.3823 0.4187 0.3994 0.4471 0.4222 0.4510 0.4857 0.4787 0.4030 0.4387 0.4442
-Long Context RAG NaiveRAG 0.3885 0.3904 0.3897 0.2267 0.2154 0.2187 0.1821 0.1475 0.1799 0.1653 0.1556 0.1649
-Propositionizer 0.1241 0.1306 0.1297 0.1074 0.1421 0.1303 0.1771 0.1645 0.1603 0.1611 0.1637 0.1656
-Compression RAG LongLLMLingua 0.3806 0.3823 0.3901 0.2155 0.2202 0.2198 0.1770 0.1567 0.1608 0.1468 0.1382 0.1493
-GraphRAG 0.3852 0.3897 0.3906 0.2213 0.2197 0.2219 0.1816 0.1770 0.1786 0.1641 0.1634 0.1625
-LightRAG 0.3930 0.3806 0.3815 0.2202 0.2216 0.2145 0.1743 0.1767 0.1799 0.1625 0.1626 0.1632
-GORAG 0.4862 0.4973 0.5192 0.4649 0.5063 0.5208 0.4814 0.4906 0.4993 0.4210 0.4340 0.4471
-Reuters NN-based Entailment 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
-Long Context RAG NaiveRAG 0.0000 0.0000 0.0375 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
-Propositionizer 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
-Compression-RAG LongLLMLingua 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0258 0.0065 0.0085
-GraphRAG 0.1375 0.1625 0.1000 0.0688 0.0813 0.0500 0.0375 0.0417 0.0417 0.0291 0.0375 0.0375
-LightRAG 0.0500 0.1375 0.1125 0.0250 0.0813 0.0563 0.0333 0.0333 0.0333 0.0125 0.0333 0.0208
-GORAG 0.0875 0.0875 0.1000 0.1750 0.1688 0.1438 0.1667 0.1958 0.2167 0.1667 0.2516 0.2193
-4.1. Experiment Settings
-4.1.1. Datasets
-We select the Web-Of-Science (WOS) dataset (Kowsari et al., 2017) and the Reuters few-shot text classification dataset by (Bao et al., 2020) for evaluating GORAG’s performance.
-We split the WOS dataset into 4, 6 and 8 rounds, and the Reuters dataset into 4 rounds, while trying to maintain a comparable number of testing data in each round.
-Due to the spacec limitation, we mainly display our experinents on the 4 rounds spit version of both datasets.
-For the WOS dataset, the texts to be classified are text chunks from academic research papers, and their labels are the respective research fields.
-In the original WOS dataset, there are 9,396 testing texts and 133 labels with label names avaiable for all rounds.
-For Reuters datasets, the texts to be classified are text chunks of news, and their labels are numbers representing the news categories.
-In the original Reuters text classification dataset, there are 310 testing texts and 31 labels for all rounds, and the label names are not avaiable.
-Further statistics of these two dataset being splitted into 4 rounds are shown in Table 2.
-Considering the differences in the properties of the two datasets, conducting experiments on both datasets provides a comprehensive and in-depth understanding of GORAG’s characteristics.
-4.1.2. Baselines
-In this paper, we compare GORAG’s performance with 7 baselines from 3 technical categories:
-NN-based Models
-  • Entailment (Xia et al., 2021): Entailment concatenates the text data with each of the label names to form multiple entailment pairs with one text sample, hence increasing the number of training data and enhance its finetuning of a RoBERTa PLM (Liu, 2019).
-Entailment then carry out classification based on these entailment pairs and convert the text classification task into a binary classification task, for detecting whether the entailment pair formed is correct.
-Long Context RAG Models
-  • NaiveRAG (Gao et al., 2023): NaiveRAG acts as a foundational baseline of current RAG models.
-When indexing, it stores text segments of the labeled texts in a vector database using text embeddings.
-When querying, NaiveRAG generates query texts’ vectorized representations to retrieve side information based on the highest similarity in their embeddings.
-  • Propositionizer (Chen et al., 2023): Propositionizer applies a fine-tuned LLM to convert side information into atomic expressions, namely propositions, to facilitate more fine-grained information retrieval than NaiveRAG.
-Compression RAG Models
-  • LongLLMLingua (Jiang et al., 2024): LongLLMLingua is a instruction aware prompt compressor model, it applies LLM’s generation perplexity to filter out un-important tokens of the model input based on the retrieved side information and the task instruction.
-  • GraphRAG (Edge et al., 2024): GraphRAG is a graph-Based RAG model that employs the LLM to extract texts’ entities and relations, which are then represented as nodes and edges in the information graph.
-GraphRAG then aggregates nodes into communities, and generates a comprehensive community report to encapsulate global information from texts.
-  • LightRAG (Guo et al., 2024): LightRAG skips the GraphRAG’s formulation of graph communities and directly retrieve nodes or pathes from created graphs.
-It calculates the embedding similarity of text extracted entities with graph nodes, to achieve a one-to-one mapping from keywords to graph nodes and retrieve all triples involved these nodes.
-According to Table 3, among these methods, Entailment achieves the state-of-the-art performance on the DFSTC task.
-For GraphRAG and NaiveRAG, we use the implementation from (nan, [n.d.]) which optimizes their original code and achieves better time efficiency while not affect the performance; For all other baselines, we use their open-sourced official implementations.
-4.1.3. Evaluation Metrics
-In this paper, we use classification accuracy as the evaluation metric.
-Given that LLMs can generate arbitrary outputs that may not precisely match the provided labels, we consider a classification correct only if the LLM’s output exactly matches the ground-truth label name or label number.
-(a) 1-shot setting.
-(b) 5-shot setting.
-Figure 4. Experiment result of WOS dataset under 1-shot and 5-shot settings.
-(a) Model Training/Indexing time on WOS dataset.
-(b) Model Training/Indexing time on Reuters dataset.
-Figure 5. Model Training/Indexing time on two datasets under 1-shot setting.
-4.1.4. Hyperparameter and Hardware Settings
+E.2. Hyperparameter and Hardware Settings
 In each round of our experiments, we handle NN and RAG-based models differently for training and indexing.
@@
 All experiments are conducted on an Intel(R) Xeon(R) Gold 5220R @ 2.20GHz CPU and a single NVIDIA A100-SXM4-40GB GPU.
-4.2. Experiment Results
-In this paper, we employ 1-shot, 5-shot, and 10-shot settings for few-shot training, where each setting corresponds to using 1, 5, and 10 labeled training samples per class, respectively.
-We omit experiments with more than 10 labeled samples per class, as for WOS dataset, there are already over 1300 labeled training data under 10-shot setting.
-4.2.1. Effectiveness Evaluation
-As shown in Table 3, GORAG achieves the best classification accuracy over all four rounds with 1-shot and 5-shot labeled indexing data on WOS dataset, it surpasses all RAG-based baselines as well as the state-of-the-art model Entailment.
-Compared with Entailment, when apply 1-shot setting, GORAG achieves at most 31.6% accuracy gain from 0.3695 to 0.4862 at the first round.
-On Reuters dataset, GORAG achieves the best classification accuracy over the last 3 rounds with 1-shot indexing data.
-Furthermore, based on the experiment results in Table 3, we can have following observations.
-Firstly, for RAG-based models other than GORAG, although they can achieve comparable or even better performance than NN-based models in the first round, they tend to suffer a more significant performance drop as the number of labels increases.
-This is because the current RAG models compress the texts and ignore the compression of the target label set.
-As a result, when the number of target labels increases, the lengthy input can make the LLM tend to give wrong classification results and suffer from the lost-in-the-middle issue.
-Secondly, compared with NN-based models, all RAG-based models other than GORAG are less sensitive to the increase of labeled data, from 1-shot to 5-shot setting on WOS dataset, the performance of Entailment increases significantly, however, for RAG-based baselines, performance tends to remain the same from 1-shot to 10-shot setting.
-This is because RAG-based models do not train any model parameters.
-Although an increase in training data can bring new information to enrich the side information, they are not always considered useful by the model and may not be retrieved.
-Consequently, the increase in labeled data may have no effect on some data samples during testing.
-On the other hand, for GORAG, the new information can introduce new paths for GORAG’s adaptive retrieval algorithm.
-This expands the influence scope of the new information from a fixed graph neighborhood to the entire information graph, making it easier for the retrieval process to benefit from the new information.
-Thirdly, for Long Context RAG models, their lengthy retrieved side information further increase the length of LLM classification inputs, and consequently make the LLM harder to detect the relevence between side information and the labels than Compression based RAG models, exacerbating the lost-in-the-middle issue and make the LLM tend to generate wrong classification results.
-4.2.2. Robustness Evaluation
-To evaluate the robustness of GORAG and several representative baselines from the three technical categories, we conducted additional experiments on the 6 and 8 round split versions of the WOS dataset under 1-shot and 5-shot settings, as the WOS dataset contains more testing data than Reuters dataset.
-In Figure 4, the results for rounds 1 to 4 are obtained from the 4-round split version, rounds 5 and 6 are obtained from the 6-round split version, and rounds 7 and 8 are obtained from the 8-round split version of the WOS dataset.
-This setting ensures that the testing data at each round is as sufficient as possible while maintaining a acceptable amount of experiment time cost.
-As illustrated in Figure 4, the classification accuracy of GraphRAG and NaiveRAG drops significantly from round 1 to round 8, demonstrating the negative impact of the lengthy and unfiltered target label set on classification results.
-However, compared to other RAG-based baselines, GORAG maintains competitive classification accuracy as the number of rounds increases through 1 to 8 round, with both of the 1-shot and 5-shot setting.
-Table 4. The size of the constructed information graph in each round.
-Dataset Model Round
-R1 R2 R3 R4 After R4
-Node # Edge # Node # Edge # Node # Edge # Node # Edge # Node # Edge #
-WOS GORAG offline 1,149 1,405 3,681 4,587 4,947 6,204 5,438 6,879 5,438 6,879
-GORAG 1,149 1,405 12,108 12,245 25,357 26,290 34,521 35,919 44,283 45,973
-Reuters GORAG offline 117 119 192 200 263 284 331 366 331 366
-GORAG 117 119 584 594 1,013 1,046 1,342 1,405 1,533 1,600
-Table 5. 0-shot experiment result on WOS dataset.
-Model Round
-R1 R2 R3 R4
-Qwen2.0-7B 0.3322 0.1790 0.1451 0.1444
-Mistral0.3-7B 0.1436 0.0540 0.0336 0.0270
-LLaMA3-8B 0.3351 0.1614 0.1161 0.0930
-GORAG 0.3305 0.2567 0.2230 0.2102
-4.2.3. Efficiency Evaluation
-To evaluate the efficiency of GORAG, we compare GORAG’s indexing time cost with other baselines’ training or indexing time cost on the 4 round split version of WOS and Reuters datasets, the results are shown in Figure 5.
-For both datasets, Propositionizer achieves the worst indexing efficiency.
-This is because it employs an additional fine-tuned LLM to convert side information into propositions.
-In this process, this extra LLM generates a long list of propositions converted from each indexing or testing text, introducing more efficiency overhead compared to other models that use the LLM solely for classification, which typically generate only a few tokens representing the label name or label number.
-On the WOS dataset, the indexing procedure of compression RAG-based models exhibits better time efficiency than the training procedure of NN-based models from R2 to R4.
-This is because NN-based models require retraining the entire model with all seen labeled data in each round, which limits their training efficiency.
-On the Reuters dataset, GraphRAG achieves the worst indexing efficiency among all Compression RAG models.
-This is because the Reuters dataset contains fewer labels than the WOS dataset in each round.
-As a result, the constructed graph tends to be smaller than that of the WOS dataset, as shown in Table 4.
-Consequently, since the time cost of GraphRAG’s community generation procedure is also determined by the number of communities generated (Traag et al., 2019), GraphRAG introduces a more significant efficiency overhead on the Reuters dataset where the created information graph is much smaller compared to the WOS dataset.
-4.3. Ablation Study
-To further study GORAG’s zero-shot ability, we conducted an experiment on the WOS dataset by providing only the label names, without any labeled data to GORAG and some widely applied open-source LLM models, the result is shown in Table 5.
-In the 0-shot setting, GORAG first generates label descriptions based on the label names, then extracts keywords from these descriptions, without using any information from labeled texts.
-Note that the Entailment model cannot be applied in this setting, as they require labeled data for training or indexing.
-For RAG-based models, their performance in the 0-shot setting is equivalent to that of their backbone LLMs due to the lack of indexing data.
-As shown by the results, GORAG achieves comparable performance with other open-source LLM models in the first round and outperforms them in all subsequent rounds.
-This is because, when the number of target labels is small, the benefit of compressing the target label set is not as pronounced.
-However, as the number of labels increases, the importance of filtering the label set becomes more significant, allowing GORAG to consistently outperform other open-source LLM models.
-To better study the reason behind GORAG’s superior performance, we conduct an extensive ablation study under 1-shot setting on GORAG.
-For the ablation studie, we only experimented with the 1-shot setting unless further illustrated, as the performance of RAG models do not change significantly with different shot settings.
-Specifically, to better study how different components of GORAG affects the classification performance, we apply its following variants for ablation study:
-  • GORAG unit: The GORAG model that remove the edge weighting mechanim, every edge in this variant is assigned with weight 1.
-  • GORAG offline: The GORAG model that remove the online indexing mechanism.
-  • GORAG keyword: The GORAG model that only use the keyword extracted from the text to create LLM classification input, rather than the whole text.
-Table 6. Ablation experiments for variants of GORAG.
-Dataset Model Round
-R1 R2 R3 R4
-WOS GORAG unit 0.4706 0.4394 0.4407 0.3899
-GORAG offline 0.3063 0.2302 0.2455 0.2156
-GORAG keyword 0.4746 0.4606 0.4455 0.4030
-GORAG 0.4862 0.4649 0.4814 0.4210
-Reuters GORAG unit 0.0000 0.0625 0.1667 0.0795
-GORAG offline 0.0000 0.1000 0.0917 0.1032
-GORAG keyword 0.0000 0.1000 0.1167 0.1438
-GORAG 0.0875 0.1750 0.1667 0.1667
+E.3. Further Analysis based on Precision and F1
+According to Table 8 and Table 2, although Entailment achieves the second best accuracy than GORAG on WOS dataset, it has a high hallucination rate.
+This is because the ground truth label of Entailment’s entailment pairs are greatly unbalanced, hence make the model easy to overfit, and tend to predict all entailment pairs of a particular text as incorrect under few-shot settings.
 (a) GORAG’s candidate selection algorithm returns candidate labels adaptively based on keywords extracted from the tesing texts.
-(b) GORAG’s candidate selection algorithm only returns the ground truth label of the testing text.
+(b) GORAG’s candidate selection algorithm only returns the ground truth label of the query text.
 Figure 6. Two case of GORAG successfully generate candidate labels to help text classification on WOS dataset.
-In Table 6, we present the results of these variants to evaluate whether the edge weighting mechanism, the online indexing mechanism, and the substitution of lengthy texts with shorter keywords can enhance the prediction accuracy of GORAG.
-The results show that GORAG achieves the best prediction accuracy, demonstrating the importance of the edge weighting and onling indexing mechanism.
-Furthermore, GORAG keyword also achieves a competitive performance, illustrating the keywords extracted from texts can also have valuable information for text classification.
+E.4. GORAG with different backbone LLMs
 Table 7. Evaluation on different LLM backbones.
@@
 GORAG Qwen2.5 0.5101 0.4839 0.4823 0.4235
-Additionally, in Table 4, we compare the size of the constructed information graph with and without online indexing after each round.
-After applying the online indexing mechanism, the information graph being significantly enriched from R1 to After R4.
 In Table 7, we present the results of GORAG with different backbone LLMs, we test GORAG with more advanced Qwen2.5-7B (Qwen Team, 2024), the result demonstates the GORAG’s performance can be further improved with more advanced LLM backbones.
-4.4. Case Study
-In this subsection, to illustrate the strength of GORAG’s adaptive retrieval and candidate label generation procedure, we dig into two testing cases select from WOS dataset where GORAG’s adaptive retrieval helps to reduce the LLM input length and benefit text classification.
+E.5. Case Study
+In this section, to illustrate the strength of GORAG’s adaptive retrieval and candidate label generation procedure, we dig into two testing cases select from WOS dataset where GORAG’s adaptive retrieval helps to reduce the LLM input length and benefit text classification.
 As shown in Figure 6(a), for the case whose ground truth label is Algorithm Design, the candidate label retrieved by GORAG are Computer programming and Algorithm design.
@@
 On the other hand, for the case in Figure 6(b) whose ground truth label is Electrical Circuits, since the ground truth label name Electrical Circuits already exists in the text and being extracted as keyword, GORAG’s adaptive retrieval algorithm is almost certain to classify this text into class Electrical Circuits, hence GORAG’s adaptive retrieval and candidate type generation algorithm would only select the only candidate label retrieved from the constructed graph and input to the LLM.
-4.5. Discussion
-According to the aforementioned experimental results, we have identified several key characteristics of NN-based models and RAG-based models.
-On the one hand, despite the rapid development of LLMs in recent years, they can still yield sub-optimal classification accuracy in DFSTC tasks due to the lengthy LLM inputs.
-When there is a sufficient amount of labeled data available in each dynamic updation round, the NN-based model Entailment can also achieve comparable or even better performance than traditional RAG based models other than GORAG.
-On the other hand, compared with NN-based models, benefited from the time efficient indexing of the side information and the rich pre-trained knowledge within the LLM, by filter out some unrelated labels and shorten the LLM inputs, RAG-based models are better suited for DFSTC task when the target labels are updated frequently and there only existes a limit number or even none of the labeled data is avaiable for each dynamic updation round.
-5. Conclusion
-In this paper, we propose GORAG, a Graph-based Online Retrieval Augmented Generation framework for the Dynamic Few-shot Text Classification (DFSTC) task.
-Extensive experiments on two text datasets with different characteristics demonstrate the effectiveness of GORAG in classifying texts with only a limited number or even no labeled data.
-Additionally, GORAG shows its effectiveness in adapting to the dynamic updates of target labels by retrieving candidate labels to filter the large target label set in each update round.
-With extensive of ablation studies, we confirm that GORAG’s adaptive retrieval, edge weighting, and online indexing mechanisms contribute to its effectiveness.
-For future work, we aim to further enhance GORAG’s performance and explore its application in more complex and diverse scenarios, as well as further improve the efficiency of GORAG’s online indexing mechanism.
+Table 8. The macro recall score and hallucination rate (Hall. Rate) on WOS dataset.
+Metrics Category Model Round
+R1 R2 R3 R4
+1-shot 5-shot 10-shot 1-shot 5-shot 10-shot 1-shot 5-shot 10-shot 1-shot 5-shot 10-shot
+Recall NN-based Entailment 0.3502 0.3475 0.3908 0.3021 0.3415 0.3442 0.2327 0.2406 0.2432 0.1789 0.1792 0.1837
+Long Context RAG NaiveRAG 0.3324 0.3417 0.3398 0.1615 0.1652 0.1650 0.0943 0.1013 0.0872 0.0568 0.0805 0.0628
+Propositionizer 0.1032 0.1131 0.1114 0.0947 0.1210 0.1187 0.1352 0.1298 0.1254 0.1373 0.1345 0.1402
+Compression RAG LongLLMLingua 0.3139 0.3105 0.3258 0.1349 0.1388 0.1350 0.0862 0.0837 0.0872 0.0616 0.0610 0.0628
+GraphRAG 0.3791 0.3646 0.3732 0.1535 0.1519 0.1557 0.1009 0.1014 0.1015 0.0902 0.0940 0.0912
+LightRAG 0.3462 0.3441 0.3519 0.1544 0.1507 0.1466 0.1241 0.1257 0.1226 0.1129 0.1180 0.1101
+GORAG 0.4387 0.4659 0.4874 0.3074 0.3472 0.3483 0.2371 0.2513 0.2445 0.1960 0.1970 0.2013
+Hal. Rate NN-based Entailment 54.53% 58.96% 51.63% 44.76% 36.62% 37.76% 32.57% 28.14% 27.35% 29.68% 26.86% 24.13%
+Long Context RAG NaiveRAG 3.00% 2.32% 2.63% 6.45% 6.16% 6.28% 6.93% 7.46% 7.52% 9.96% 9.50% 9.83%
+Propositionizer 6.85% 5.88% 6.97% 7.40% 6.52% 6.09% 5.97% 6.44% 6.30% 5.78% 6.11% 5.52%
+Compression-RAG LongLLMLingua 6.50% 6.29% 5.63% 16.45% 16.05% 16.28% 19.63% 20.47% 20.52% 21.96% 21.35% 21.83%
+GraphRAG 2.40% 2.23% 2.07% 6.03% 5.97% 6.18% 7.20% 7.58% 7.03% 6.90% 6.39% 6.49%
+LightRAG 1.70% 1.24% 1.37% 2.59% 2.78% 3.21% 4.66% 4.55% 4.58% 5.89% 5.80% 5.73%
+GORAG 0.66% 1.16% 0.74% 0.30% 0.68% 0.65% 0.41% 0.57% 0.68% 0.40% 0.43% 0.64%
+Table 9. The macro recall score and hallucination rate (Hall. Rate) on IFS-Rel dataset.
+Metrics Category Model Round
+R1 R2 R3 R4
+1-shot 5-shot 10-shot 1-shot 5-shot 10-shot 1-shot 5-shot 10-shot 1-shot 5-shot 10-shot
+Recall NN-based Entailment 0.3191 0.6015 0.5515 0.2172 0.3043 0.3330 0.1225 0.1802 0.2081 0.1093 0.1267 0.1490
+Long Context RAG NaiveRAG 0.4039 0.4194 0.4292 0.2320 0.2009 0.2115 0.1147 0.1163 0.1097 0.0871 0.0895 0.0912
+Propositionizer 0.1053 0.1146 0.1104 0.1124 0.1201 0.1197 0.0824 0.1029 0.0902 0.0744 0.0717 0.0799
+Compression RAG LongLLMLingua 0.4426 0.4625 0.4044 0.1879 0.1559 0.2158 0.0947 0.0906 0.1113 0.0698 0.0822 0.0696
+GraphRAG 0.4264 0.4139 0.4208 0.2116 0.2008 0.2047 0.1213 0.1179 0.1185 0.0906 0.0857 0.0883
+LightRAG 0.4221 0.3985 0.4382 0.2344 0.2150 0.2429 0.1270 0.1313 0.1276 0.0902 0.0926 0.0918
+GORAG 0.4588 0.4681 0.5000 0.2348 0.2175 0.2405 0.1279 0.1295 0.1290 0.0950 0.0886 0.0919
+Hal. Rate NN-based Entailment 8.10% 15.63% 4.69% 26.41% 19.14% 6.25% 26.56% 17.86% 6.67% 6.48% 5.90% 3.28%
+Long Context RAG NaiveRAG 1.56% 4.25% 1.78% 3.05% 2.97% 3.13% 3.02% 3.70% 3.54% 4.02% 4.18% 3.63%
+Propositionizer 7.82% 7.34% 11.25% 5.78% 6.40% 6.72% 4.84% 6.09% 7.03% 8.13% 8.59% 6.72%
+Compression-RAG LongLLMLingua 3.59% 2.97% 6.56% 2.12% 6.23% 5.47% 6.36% 6.00% 3.50% 3.29% 5.65% 8.88%
+GraphRAG 4.25% 1.56% 1.25% 2.66% 3.98% 2.89% 3.39% 3.75% 3.80% 3.91% 3.24% 3.09%
+LightRAG 3.28% 2.66% 2.03% 2.73% 2.97% 2.27% 1.98% 1.56% 1.67% 2.46% 2.23% 2.30%
+GORAG 0.31% 0.63% 0.16% 1.78% 2.81% 2.58% 1.30% 0.78% 1.09% 1.48% 1.02% 1.25%